### v1

# Pricing and Hedging Strategies for Cross-Currency Equity Protection Swaps

Abstract
In this paper, we explore the pricing and hedging strategies for an innovative insurance product called the equity protection swap (EPS). Notably, we focus on the application of EPSs involving cross-currency reference portfolios, reflecting the realities of investor asset diversification across different economies. The research examines key considerations regarding exchange rate fluctuations, pricing and hedging frameworks, in order to satisfy dynamic requirements from EPS buyers. We differentiate between two hedging paradigms: one where domestic and foreign equities are treated separately using two EPS products and another that integrates total returns across currencies. Through detailed analysis, we propose various hedging strategies with consideration of different types of returns — nominal, effective, and quanto - for EPS products in both separate and aggregated contexts. The aggregated hedging portfolios contain basket options with cross-currency underlying asset, which only exists in the OTC market, thus we further consider a superhedging strategy using single asset European options for aggregated returns. A numerical study assesses hedging costs and performance metrics associated with these hedging strategies, illuminating practical implications for EPS providers and investors engaged in international markets. We further employ Monte Carlo simulations for the basket option pricing, together with two other approximation methods - geometric averaging and moment matching. This work contributes to enhancing fair pricing mechanisms and risk management strategies in the evolving landscape of cross-currency financial derivatives.
1 Introduction
Xu et al. [19] have introduced a new type of insurance for superannuation - equity protection swap (EPS). It is a financial derivative, which is reminiscent of a total return swap but has also some features of the annuity insurance product RILA. Additionally, the structure of an EPS makes it a reliable safeguard when the value of a reference portfolio drops, and it can also be tailored to accommodate the specific requirements of super fund members. Compared with variable annuity riders, an EPS is less complex and is not necessarily associated with a variable annuity. Apart from that, there is a main difference between RILA and EPS products that EPS is an insurance product that can be added to the variable annuity rather than the annuity itself in order to manage the risks.
An important direction in the study of EPSs is the case of a cross-currency reference portfolio, which will be discussed in this paper. In reality, it is fairly common for investors to hold assets across different economies in order to do risk diversification. Then the underlying reference portfolio for an EPS will contain assets from several economies. Therefore, to generalise the application scope of the concept of an EPS, the pricing and hedging strategies should take into account the exchange rate fluctuations, as well as the correlation between prices of domestic and foreign assets. Various existing approaches to the arbitrage-free pricing and hedging of cross-currency reference portfolios include either considering two independent local returns on domestic and foreign assets or focusing on the total return on the portfolio’s value expressed in the domestic currency. This area of research is essential for providers and investors operating in international markets and it provides a valuable insight into fair pricing and hedging strategies for an EPS in a global context.
To mitigate the foreign exchange risk, a variant of the Black-Scholes model was first proposed for pricing and hedging of European currency options by Garman and Kohlhagen [7]. Since the Garman-Kohlhagen model is based on the geometric Brownian motion, it is unable to correctly specify the dynamics of the exchange rate. Later, Shokrollahi and Kılıçman [18] proposed the actuarial approach for pricing currency options with exchange rate governed by the jump process and mixed fractional Brownian motion by changing the currency options pricing problem into an equivalent fair insurance premium problem.
In order to hedge the risk of investments in domestic and foreign reference portfolios, we need to consider the preferences of EPS buyers and examine their respective strategies. If the purchaser regards the underlying portfolio as two separate independent portfolios, they will focus on the actual returns of domestic and foreign equities independently. In this case, we can model two hedging portfolios to hedge domestic and foreign equity risks separately. We need to mention that the hedging portfolio for domestic equity has been discussed before, thus we only need to consider foreign equity risks in this strategy. In another situation, EPS buyers only focus on the total return on the whole reference portfolio rather than separate returns of domestic and foreign equities. The insurance company should hedge total return risks together with currency risks and build only one hedging portfolio, it is more difficult to structure than in the first case.
We aim to analyse the pricing and hedging of an EPS with a cross-currency reference portfolio. For conciseness of notation, we assume that the reference portfolio contains only two economies - a domestic and a foreign one. By convention, the domestic currency is chosen to be the Australian dollar (AUD) and the foreign one is the U.S. dollar (USD) but, obviously, all definitions and results from this paper can be applied to any two (or more than two) currencies.
We start this paper by defining a setup for the domestic and foreign markets in Section 2. We discuss not only the nominal return but also the effective return of the foreign equity, then provide the aggregated effective and quanto (i.e., for a guaranteed exchange rate) rate of return for the cross-currency portfolio. The cross-currency reference portfolio of EPS is defined and we discuss the hedging strategies for such EPS products. Unlike for a standard EPS with a single currency reference portfolio studied by Xu et al. [19], here we have several different hedging strategies since we consider alternative conventions regarding realised returns and hence various structures of the protection and fee legs.
In Section 3 , we study separate hedging strategies, which correspond to the EPS convention where the realised returns of domestic and foreign equities are addressed in separation and, as a consequence, two sets of hedging portfolios with domestic and foreign options are used. In fact, in that case we can adapt the static hedging strategies for a standard EPS with a single currency reference portfolio to fully hedge the domestic equity, by adding appropriate weights to the notional principal. However, for the foreign equity, we need to consider not only its realised return but also the returns on the exchange rate. Firstly, the EPS provider can hedge only the nominal return of the foreign equity and exchange this foreign nominal return to domestic return by multiplying the exchange rate at maturity. In this case, European options written on foreign equities can be used directly to build the hedging portfolio and the volatility of the exchange rate will not be considered.
Secondly, one can also hedge the foreign equity’s effective return, which is the return on the foreign equity struck in domestic currency. The EPS provider needs to track the value of foreign equity expressed in the domestic currency in the holding period of the EPS product and, in this case, he needs to use European options with foreign equity struck in the domestic currency as the underlying asset. In the second separate hedging strategy, all of the values and realised returns are settled in the domestic currency. Besides, we also consider hedging the foreign equity’s quanto rate in Section 3.5, which predetermines a fixed exchange rate corresponding to the EPS holder’s requirements.
In Section 4, we consider the hedging strategy for an EPS referencing the total return that combines the returns of both domestic and foreign equities. We discuss two different hedging strategies here, the first one is a static hedging strategy that uses European cross-currency basket options with the whole portfolio struck in domestic currency as the underlying asset; here we can extend the algorithm of the hedging strategy for standard EPS products with single currency reference portfolio to do the hedging. However, we need to mention that this fully hedging strategy may be unattainable in that European cross-currency basket options on bespoke portfolios are unlikely to be offered in practice, though they can be negotiated on the OTC market. In order to build a more practical hedging portfolio, we further introduce a joint superhedging strategy. By using the subadditive property, we build a joint superhedging strategy by using European options with single currency underlying assets. However, some parts of this joint superhedging strategy still need to be structured in the OTC market.
In Section 5, we present a numerical study for EPS products with the cross-currency reference portfolio. We select parameters for domestic and foreign markets and examine trends of stock prices and exchange rates with correlated volatility. Our numerical study illustrated the hedging costs for separate hedging strategies: domestic return, nominal foreign return, effective foreign return, and quanto foreign return, as well as for aggregated hedging strategies: effective hedging strategies, quanto hedging strategies, and superhedging strategies. For the aggregated hedging strategies, we need to use basket options to build the hedging portfolio. However, we cannot find an explicit formula for the basket option and we will use some approximations instead. We will consider the Monte Carlo simulation as an ”actual” cost, the closed-form solutions for basket options with geometric averaging approximation and under three-moment matching approximation method are also included as a comparison.
Each hedging strategy satisfies different requirements from EPS buyers, which influence the structure and clauses of the EPS product at the beginning, thus all of them are important and cannot be substituted. In the meantime, the aggregated superhedging strategy requires higher hedging costs compared with other aggregated hedging strategies, it is consistent with the definition of superhedging. We conclude this work by summarising our findings and emphasising the practical importance of this study.
2 Equity Protection Swaps
In this section, we define a generic equity protection swap (EPS), which is a financial contract with a single terminal payoff at maturity $T$ determined by the nominal principal $N$ and the simple rate of return of an underlying reference portfolio over the lifetime $[0,T]$ of an EPS. We denote by $R_{T}$ the simple rate of return on the strictly positive wealth $X$ of a reference portfolio so that $R_{T}:=(X_{T}-X_{0})/X_{0}$ satisfies $R_{T}>-1$. Observe that in the cross-currency framework the wealth process $X$ can be defined in several alternative ways, depending on the interpretation of a portfolio $X$ and investor’s preferences. According to the definition of standard EPS, which was introduced in Xu et al. [19], the cash flow at time $T$ from the investor to the provider of an EPS is given by $N\psi(R_{T})$ where the function $\psi:(-1,\infty)\to{\mathbb{R}}$ is used to encode the structure of an EPS. We postulate that the payoff profile $\psi$ is a piece-wise linear, non-decreasing, continuous function such that $\psi(0)=0$. We first formulate the definition of a generic EPS. Unless explicitly stated otherwise, we assume, without loss of generality, that the notional principal of an EPS is one unit of the domestic currency and all results can be scaled by the actual value of the nominal principal. The following definition was introduced in Xu et al. [19].
Definition 2.1.
In a standard equity protection swaps starting at time 00 with the maturity date $T$ and the reference portfolio $X$, the cash flow for the EPS provider is specified by the adjusted return $\psi(R_{T})$ where the derivative of the adjusted return function $\psi:\mathbb{R}\to\mathbb{R}$ equals
$\psi^{\prime}(R)=\sum_{k=-(n+1)}^{m}\alpha_{k+1}\mathds{1}_{(\beta_{k},\beta_{k+1})}(R)$ (1)
where the sequence of endpoints of return ranges, $(\beta_{k})_{k=-n}^{m}$, is a strictly increasing finite sequence of real numbers, $\beta_{-(n+1)}=-1$ and $\beta_{m+1}=\infty$. The sequence $(\alpha_{k})_{k=-n}^{m+1}$ of marginal participation rates is an arbitrary sequence of numbers from $[0,1]$. Furthermore, we assume that $\beta_{0}=0$ and thus $-1=\beta_{-(n+1)}<\beta_{-n}<\cdots<\beta_{-2}<\beta_{-1}<\beta_{0}=0<\beta_{1}<\cdots<\beta_{m}<\beta_{m+1}=\infty$.
For simplicity, we will write $-1=l_{n+1}<l_{n}<\cdots<l_{2}<l_{1}<l_{0}=0$ for non-positive values of $\beta$ and $0=g_{0}<g_{1}<\cdots<g_{m}<g_{m+1}=\infty$ for non-negative values of $\beta$. For each $k$, the protection rate over the loss interval $[l_{k+1},l_{k})$ is denoted as $p_{k+1}$ whereas the fee rate over the gain interval $(g_{k},g_{k+1}]$ is denoted as $f_{k+1}$. Then we have the following lemma (see Xu et al. [19]).
Lemma 2.1.
We have that $\psi(R)\leq 0$ for $R\in(-1,0]$ and $\psi(R)\geq 0$ for $R\in[0,\infty)$. For $i=0,1,\dots,n$ and $j=0,1,\dots,m$, the protection leg and the fee leg of an EPS are given by
$\psi^{p}(R)$ $:=\psi(R)\mathds{1}_{\{R<0\}}=\Big{(}\sum_{k=0}^{i-1}p_{k+1}\Delta l_{k}+p_{i+1}(R-l_{i})\Big{)}\mathds{1}_{\{R\in(l_{i+1},l_{i})\}},$
$\psi^{f}(R)$ $:=\psi(R)\mathds{1}_{\{R>0\}}=\Big{(}\sum_{k=0}^{j-1}f_{k+1}\Delta g_{k}+f_{j+1}(R-g_{j})\Big{)}\mathds{1}_{\{R\in(g_{j},g_{j+1})\}},$
where $\Delta l_{i}=l_{i+1}-l_{i}<0$ and $\Delta g_{j}=g_{j+1}-g_{j}>0$. We have $\psi(R_{T})=\psi^{p}(R_{T})+\psi^{f}(R_{T})$, where the non-positive payoff $\psi^{p}(R_{T})$ and the non-negative payoff $\psi^{f}(R_{T})$ represent the provider’s loss (protection payout) and gain (fee income) from an EPS, respectively.
2.1 Buffer and Floor EPS
Two standard EPS products were introduced by Xu et al. [19] and will be used here as examples of a standard EPS in further analysis. Needless to say, much more complex products can be offered to cater for individual risk preferences of investors and all of them can be dealt with using general pricing and hedging results from this work. However, for simplicity, we only deal here with EPS products of European style with a payoff at maturity date $T$, though one may allow for cancellation of an EPS by its holder at its market price if the underlying European options are liquidly traded.
We first define the buffer EPS where the concept of a buffer applies to both the protection and fee legs and thus it could also be called an EPS with a double buffer.
Definition 2.2.
The buffer EPS is obtained by taking $p_{1}=f_{1}=0$ (hence $\psi_{B}(R)=0$ for all $R\in[l_{1},g_{1}]$ where $l_{1}<0$ and $g_{1}>0$) and some values for $p_{2}\in(0,1]$ and $f_{2}>0$. Hence the cash flow at maturity $T$ of a buffer EPS to its provider equals
$\psi_{B}(R_{T})$ $:=p_{2}(R_{T}-l_{1})\mathds{1}_{\{R_{T}\in(-1,l_{1})\}}+f_{2}(R_{T}-g_{1})\mathds{1}_{\{R_{T}\in(g_{1},\infty)\}}$
$=-p_{2}(l_{1}-R_{T})^{+}+f_{2}(R_{T}-g_{1})^{+}$
$=\frac{p_{2}}{X_{0}}\,\text{Put}_{0}(K^{l}_{1}(X_{0}),T)-\frac{f_{2}}{X_{0}}\,\text{Call}_{0}(K^{g}_{1}(X_{0}),T)$
where $K^{l}_{1}(X_{0})=(1+l_{1})X_{0}$ and $K^{g}_{1}(X_{0})=(1+g_{1})X_{0}$.
In the floor EPS we set $p_{2}=0,p_{1}\in(0,1],f_{1}=0,f_{2}\in(0,1]$, which means that the provider covers the holder’s losses above the level $l_{1}$ but the holder’s losses below $l_{1}$ are not covered at all. Notice that, by convention, the buffer is applied to the fee leg and thus it is exactly the same as in a buffer EPS.
Definition 2.3.
The floor EPS is specified by $p_{2}=0,p_{1}\in(0,1],f_{1}=0,f_{2}>0$ and thus the cash flow at maturity $T$ to its provider equals
$\psi_{F}(R_{T})$ $:=p_{1}l_{1}\mathds{1}_{\{R_{T}\in(-1,l_{1}]\}}-p_{1}(-R_{T})^{+}\mathds{1}_{\{R_{T}\in(l_{1},0]\}}+f_{2}(R_{T}-g_{1})^{+}$
$=-p_{1}(-R_{T})^{+}+p_{1}(l_{1}-R_{T})^{+}+f_{2}(R_{T}-g_{1})^{+}$
$=-\frac{p_{1}}{X_{0}}\,\text{Put}_{T}(K^{l}_{1}(X_{0}),T)+\frac{p_{1}}{X_{0}}\,\text{Put}_{T}(X_{0},T)-\frac{f_{2}}{X_{0}}\,\text{Call}_{T}(K^{g}_{1}(X_{0}),T).$
2.2 Static Hedging and Fair Pricing of an EPS
After representing the payoff of a generic EPS, one can observe that its payoff can be written in terms of a combination of European calls and puts on rate of return $R_{T}$ maturing at $T$. Hence our next goal is to examine a hedging portfolio for a generic EPS in terms of options written on the underlying portfolio $X$.
Definition 2.4.
A static portfolio of European put and call options is represented by a vector $h=(a_{0},a_{1},a_{2},\dots,a_{n},b_{0},b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{n+m}$ and its value process $H$ is given by, for every $t\in[0,T]$,
$H_{t}:=\sum_{i=0}^{n}a_{i}\,\text{\rm Put}_{t}(X,K^{l}_{i}(X_{0}),T)+\sum_{j=0}^{m}b_{j}\,\text{\rm Call}_{t}(X,K^{g}_{j}(X_{0}),T)$
where for $t=T$ we have that $\text{\rm Put}_{\,T}(X,K^{l}_{i}(X_{0}),T)=(K^{l}_{i}(X_{0})-X_{T})^{+}$ where the strike equals $K^{l}_{i}(X_{0})=(1+l_{i})X_{0}$ and $\text{\rm Call}_{\,T}(X,K^{g}_{j}(X_{0}),T)=(X_{T}-K^{g}_{j}(X_{0}))^{+}$ where $K^{g}_{j}(X_{0})=(1+g_{j})X_{0}$.
According to our convention, the provider receives from the buyer the initial premium $c\in\mathbb{R}$ at time 0 and the cash flow $\psi(R_{T})$ at time $T$. Assume that the provider establishes at time 0 a portfolio $h$ of plain-vanilla call and put options written on $X$. Then we have the following definition for the provider’s hedged cash flow and fair price.
Definition 2.5.
For an EPS introduced in Definition 2.1, the provider’s hedged cash flow at time $T$ per one unit of the nominal principal $N$, denoted by $\mbox{\rm CF$\,$}_{T}(c,h)$, equals
$\mbox{\rm CF$\,$}_{T}(c,h)=(c-H_{0})e^{rT}+H_{T}+\psi(R_{T}).$
We say that an EPS can be statically hedged by the provider if there exists a static portfolio $\widehat{h}$ established at time 0 and an initial premium $\widehat{c}\in\mathbb{R}$ such the equality $\mbox{\rm CF$\,$}_{T}(\widehat{c},\widehat{h})=0$ holds almost surely. Then the number $\widehat{c}$ is called the fair premium for an EPS per one unit of the nominal principal.
Notice that it suffices to find $\widehat{h}$ such that the $\widehat{H}_{T}=N\psi(R_{T})$ where $N$ is the nominal principal of an EPS and set $\widehat{c}=\widehat{H}_{0}$. The static hedge and fair price for a standard EPS with arbitrary payoff structure can be identified from the following proposition, which is due to Xu et al. [19] (see Proposition 3.1 therein). Recall that $l_{0}=g_{0}=0$ and, for convenience, we also adopt the convention that $f_{0}=p_{0}=0$.
Proposition 2.1.
The static hedge $(\widehat{c},\widehat{h})$ composed of call and put options for a standard EPS with the nominal principal $N$ satisfies the following equation
$\widehat{H}_{T}=N\sum_{i=0}^{n}\frac{p_{i+1}-p_{i}}{X_{0}}\,\mbox{\rm Put}_{T}(X,K^{l}_{i}(X_{0}),T)-N\sum_{j=0}^{m}\frac{f_{j+1}-f_{j}}{X_{0}}\,\mbox{\rm Call}_{T}(X,K^{g}_{j}(X_{0}),T)$ (2)
where $K^{l}_{i}(X_{0})=(1+l_{i})X_{0}$ and $K^{g}_{j}(X_{0})=(1+g_{j})X_{0}$ for every $i=0,1,\dots,n$ and $j=0,1,\dots,m$. The fair premium $\widehat{c}$ for an EPS is equal to $\widehat{H}_{0}$ per one unit of the nominal principal where, in the domestic currency,
$\widehat{H}_{0}=\sum_{i=0}^{n}\frac{p_{i+1}-p_{i}}{X_{0}}\,\mbox{\rm Put}_{0}(X,K^{l}_{i}(X_{0}),T)-\sum_{j=0}^{m}\frac{f_{j+1}-f_{j}}{X_{0}}\,\mbox{\rm Call}_{0}(X,K^{g}_{j}(X_{0}),T).$ (3)
The proof of Proposition 2.1 hinges on the following elementary equalities, for every $i=0,1,\dots,n$,
$(l_{i}-R_{T})^{+}=\frac{((1+l_{i})X_{0}-X_{T})^{+}}{X_{0}}=\frac{1}{X_{0}}\,\mbox{\rm Put}_{T}(X,(1+l_{i})X_{0},T)$
and, for every $j=0,1,\dots,m$,
$(R_{T}-g_{j})^{+}=\frac{(X_{T}-(1+g_{j})X_{0})^{+}}{X_{0}}=\frac{1}{X_{0}}\,\mbox{\rm Call}_{T}(X,(1+g_{j})X_{0},T).$
Remark 2.1.
Let us define $\widehat{X}_{t}:=(X_{0})^{-1}X_{t}$ so that $\widehat{X}_{0}=1$ and the rates of return on processes $\widehat{X}$ and $X$ manifestly coincide. Then we obtain an equivalent representation for equality (2)
$\widehat{H}_{T}=N\sum_{i=0}^{n}(p_{i+1}-p_{i})\,\mbox{\rm Put}_{T}(\widehat{X},1+l_{i},T)-N\sum_{j=0}^{m}(f_{j+1}-f_{j})\,\mbox{\rm Call}_{T}(\widehat{X},1+g_{j},T),$ (4)
which shows that, formally, it suffices to examine options written on the normalized process $\widehat{X}$. In Section 3, the process $X$ is assumed to represent a market quantity (e.g., the price of a given equity or a specific domestic or foreign market index) we will use representation (2), which is based on traded options written on $X$.
In contrast, in Section 4 we prefer to employ equality (4) since, firstly, options on a bespoke cross-currency portfolio $X$ are unlikely to be traded and, secondly, equation (4) yields simpler pricing and hedging formulae within the cross-currency market model of Section 3.2. To conclude, equality (2) is used when the emphasis is put on the possibility of static hedging of an EPS with traded options on market indices, whereas (4) is preferred when performing computations within a stochastic market model, especially when dealing with basket options on bespoke portfolios composed of domestic and foreign assets.
Remark 2.2.
We say that an EPS is fair if its fair premium vanishes, that is, $\widehat{c}=0$ in Definition 2.5. In the case of a buffer or floor EPS, assuming that all parameters except for the fee rate $f_{2}$ are fixed, one can find a unique level of $f_{2}$ for which the fair premium for an EPS is null by solving for $f_{2}$ the linear equation (3) (or, equivalently, (4)) with initial hedging cost $\widehat{H}_{0}$ set to 0. For instance, the fair fee rate for a buffer EPS equals
$\widehat{f}_{2}=p_{2}\,\frac{\mbox{\rm Put}_{0}(X,K^{l}_{1}(X_{0}),T)}{\mbox{\rm Call}_{0}(X,K^{g}_{1}(X_{0}),T)}=p_{2}\,\frac{\mbox{\rm Put}_{0}(\widehat{X},1+l_{1},T)}{\mbox{\rm Call}_{0}(\widehat{X},1+g_{1},T)}$
and for the floor EPS we have
$\widehat{f}_{2}=p_{1}\,\frac{\mbox{\rm Put}_{0}(X,K^{l}_{1}(X_{0}),T)-\mbox{\rm Put}_{0}(X,X_{0},T)}{\mbox{\rm Call}_{0}(X,K^{g}_{1}(X_{0}),T)}=p_{1}\,\frac{\mbox{\rm Put}_{0}(\widehat{X},1+l_{1},T)-\mbox{\rm Put}_{0}(\widehat{X},1,T)}{\mbox{\rm Call}_{0}(\widehat{X},1+g_{1},T)}.$
If more than one nonzero fee rate appears in the specification of an EPS, then the question of its fairness may have several solutions but it is always easy to structure an EPS in such a way that its fairness is ensured.
3 Equity Protection Swaps on Domestic and Foreign Market Indices
An investor has two options: either to obtain a protection against large losses from their domestic and foreign investments using two separate EPSs, one for each reference portfolio, or to enter into a single EPS furnishing protection a total loss from the holder’s combined portfolio with value expressed in the domestic currency through a particular method. We first assume that the investor decides to purchase two independent products yielding protection against negative returns on $S^{d}$ and negative returns on $S^{f}$. Then it suffices to examine two separate hedging portfolios, the first one for the domestic component of holder’s portfolio the with notional principal $N^{d}$ in AUD and the second one for the foreign component with the notional principal $N^{f}$ in USD (or, equivalently, $N^{f,d}$ in AUD). We assume that the domestic EPS with the profile $\psi_{d}$ (resp. the foreign EPS with the profile $\psi_{f}$) is applied to the domestic return $R^{d}_{T}$ (resp. to the effective foreign return $R^{f,e}_{T}$).
3.1 Cross-Currency Portfolios
Let $S^{d}_{t}$ and $S^{f}_{t}$ be the values at time $t$ of domestic and foreign reference portfolios (typically, domestic and foreign market indices) with values expressed in the domestic and foreign currency, respectively. Unless explicitly stated otherwise, we present a model-free approach, which means that we do not use any particular stochastic processes to model the dynamics of value processes $S^{d}$ and $S^{f}$. The same remark applies to the exchange rate process, which is denoted as $Q$. The process $Q$ represents the direct quote, which means that $Q_{t}$ gives the price at time $t$ of one unit of the foreign currency (by convention, the American dollar) in terms of variable number of units of the domestic currency (the Australian dollar).
We introduce the following notation for the nominal domestic rate of return (or, simply, the nominal domestic return) and the nominal foreign rate of return, as well as the nominal return on the foreign currency, for every $t\in[0,T]$,
$R^{d}_{t}:=\frac{S^{d}_{t}-S^{d}_{0}}{S^{d}_{0}},\quad R^{f}_{t}=\frac{S^{f}_{t}-S^{f}_{0}}{S^{f}_{0}},\quad R^{q}_{t}=\frac{Q_{t}-Q_{0}}{Q_{0}}.$
Another important concept is the effective foreign rate of return (for brevity, the effective foreign return), that is, the rate of return on the foreign reference portfolio $S^{f}$ when its value at any date $t$ is expressed in the domestic currency using the current exchange rate $Q_{t}$. The effective foreign return satisfies, for every $t\in[0,T]$,
$R^{f,e}_{t}:=\frac{S^{f,e}_{t}-S^{f,e}_{0}}{S^{f,e}_{0}}=\frac{Q_{t}S^{f}_{t}-Q_{0}S^{f}_{0}}{Q_{0}S^{f}_{0}}=R^{f}_{t}+R^{q}_{t}+R^{f}_{t}R^{q}_{t}$
where $S^{f,e}_{t}:=Q_{t}S^{f}_{t}$. Notice that the nominal domestic rate of return can also be called the effective domestic rate of return from the point of view of a domestic investor who is invariably assessing his gains and losses in the domestic currency.
Let us introduce the notation for the notional principal of the holder’s cross-currency portfolio and for its partition into the domestic and foreign holdings. We consider an agent who invests at time 0 in a portfolio of domestic and foreign risky assets with a total wealth equal to $N$ (in AUD). Then the notional principal $N$ can be split into holdings of domestic and foreign assets as follows $N=N^{d}+N^{f,d}$ where $N^{d}$ (in AUD) is the notional principal of domestic assets in holder’s portfolio and $N^{f,d}$ (also given in AUD) is the notional principal of their foreign assets expressed in the domestic currency. We also denote by $N^{f}$ the notional principal of foreign assets in holder’s portfolio expressed in the foreign currency, which means that $N^{f}:=(Q_{0})^{-1}N^{f,d}$ (in USD) and thus we obtain an equivalent representation $N=N^{d}+Q_{0}N^{f}$. We find it convenient to introduce the weight $w:=N^{d}/N$, which gives the proportion of the total wealth invested in domestic assets and the weight $1-w:=N^{f,d}/N$ representing the proportion of the total wealth invested in foreign assets. For simplicity of notation, in what follows we use $N$ as the notional principal of both the holder’s portfolio and a related EPS.
It is useful to introduce notation for the number of units of the domestic and foreign reference portfolios $S^{d}$ and $S^{f}$ held in a cross-currency portfolio by setting
$N^{d}=wN=\alpha_{0}S^{d}_{0},\quad N^{f,d}=(1-w)N=\beta_{0}S^{f,e}_{0}$
where the constants $\alpha_{0}$ and $\beta_{0}$, which represent holdings in $S^{d}$ and $S^{f}$, respectively, satisfy
$\alpha_{0}:=\frac{N^{d}}{S^{d}_{0}}=\frac{wN}{S^{d}_{0}},\quad\beta_{0}:=\frac{N^{f,d}}{S^{f,e}_{0}}=\frac{(1-w)N}{S^{f,e}_{0}}=\frac{(1-w)N}{Q_{0}S^{f}_{0}}.$
Then the cross-currency portfolio in domestic asset $S^{d}$ and foreign asset $S^{f}$ with the notional principal $N$ and weight $w$ can be represented by the pair $(\alpha_{0},\beta_{0})$.
3.2 Cross-Currency Market Model
Before we analyse hedging and arbitrage-free pricing for EPSs with the cross-currency portfolio as a reference portfolio, let us briefly summarize the classical cross-currency model as presented, for instance, in Chapter 13 of Musiela and Rutkowski [16]. The domestic and foreign bank accounts are given by $dB_{t}^{d}=r^{d}B_{t}^{d}\,dt$ and $dB_{t}^{f}=r^{f}B_{t}^{f}\,dt$ for some constants $r^{d}$ and $r^{f}$. For a fixed horizon date $T>0$, the dynamics of domestic and foreign equities, $S^{d}$ and $S^{f}$, and the exchange rate, $Q$, under the physical probability measure ${\mathbb{P}}$ are specified by the SDEs
$dS^{d}_{t}=S^{d}_{t}\big{(}\mu_{t}^{d}\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}\big{)},$
$dS^{f}_{t}=S^{f}_{t}\big{(}\mu_{t}^{f}\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}\big{)},$
$dQ_{t}=Q_{t}\big{(}\mu_{t}^{q}\,dt+\bm{\sigma}^{q}\,d\textbf{W}_{t}\big{)},$
with strictly positive initial conditions $S^{d}_{0},S^{f}_{0}$ and $Q_{0}$ where $\textbf{W}=[W^{1},W^{2},W^{3}]^{*}$ where denotes the transpose is a standard Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ and the volatilities are given given by: $\bm{\sigma}^{d}=[\sigma^{1d},\sigma^{2d},\sigma^{3d}],\,\bm{\sigma}^{f}=[\sigma^{1f},\sigma^{2f},\sigma^{3f}]$ and $\bm{\sigma}^{q}=[\sigma^{1q},\sigma^{2q},\sigma^{3q}]$. Then the dynamics of the domestic and foreign equity and the exchange rate under an equivalent probability measure ${\mathbb{Q}}^{d}$, which is the domestic martingale measure, are given as
$dS^{d}_{t}=S^{d}_{t}\big{(}r^{d}\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}\big{)},$
$dS^{f}_{t}=S^{f}_{t}\big{(}(r^{f}-\bm{\sigma}^{f}(\bm{\sigma}^{q})^{*})\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}^{d}\big{)},$ (5)
$dQ_{t}=Q_{t}\big{(}(r^{d}-r^{f})\,dt+\bm{\sigma}^{q}\,d\textbf{W}_{t}^{d}\big{)},$
where $\textbf{W}^{d}$ is a standard Brownian motion on $(\Omega,\mathcal{F},{\mathbb{Q}}^{d})$. Finally, the process $\textbf{W}_{t}^{f}=\textbf{W}_{t}^{d}-(\bm{\sigma}^{q})^{*}\,t$ is a standard Brownian motion on $(\Omega,\mathcal{F},{\mathbb{Q}}^{f})$ where ${\mathbb{Q}}^{f}$ is the foreign martingale measure and thus the processes $S^{d},S^{f}$ and $Q$ are governed under ${\mathbb{Q}}^{f}$ by the following SDEs
$dS^{d}_{t}=S^{d}_{t}\big{(}(r^{d}+\bm{\sigma}^{d}(\bm{\sigma}^{q})^{*})\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}^{f}\big{)},$
$dS^{f}_{t}=S^{f}_{t}\big{(}r^{f}\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}^{f}\big{)},$
$dQ_{t}=Q_{t}\big{(}(r^{d}-r^{f}+\bm{\sigma}^{q}(\bm{\sigma}^{q})^{*})\,dt+\bm{\sigma}^{q}\,d\textbf{W}_{t}^{f}\big{)}.$
If the volatility matrix $\Sigma=[\bm{\sigma}^{d*},\bm{\sigma}^{f*},\bm{\sigma}^{q*}]$ is non-singular, then the cross-currency market model given by (3.2) is arbitrage-free and complete.
The model parameters used in our numerical studies presented in Section 5 and examples are reported in Table 1. The prices of plain-vanilla domestic or foreign options are given by the Black-Scholes formula and the pricing formulae for other options of our interest are given in Propositions 3.1, 3.2, 4.1, 4.2, 4.3, 4.4, 5.1, 5.2, 5.3 and 5.4
It is worth noting that initial prices of the foreign and domestic equity is immaterial when pricing equity protection swaps within the setup of a cross-currency market model. This is easy to understand since the payoffs of equity protection swaps only depend on realized returns on the reference domestic and foreign portfolios, as well as on respective notional principals, but they do not depend on the initial prices of equities. In contrast, when presenting static hedging portfolios on various kinds of EPSs, we assume that the trading strategies are static and based on actively traded options and thus the initial prices of equities and real-world strikes of traded options are relevant.
To sum up, one should note the distinction between the static hedging results based on traded options and theoretical computations where prices are computed within a cross-currency market model using an implicit assumption that all options can be dynamically hedged using underlying domestic and foreign assets.
3.3 Protection Against Domestic Losses
To protect the domestic component of his portfolio with the initial wealth $N$ and weight $w$, the holder may use the domestic EPS with the notional principal $N^{d}=wN$ and the payoff at time $T$ given in the domestic currency
$\mbox{\rm EPS\,}^{d}_{T}:=N^{d}\psi_{d}(R^{d}_{T})=wN\psi_{d}(R^{d}_{T}).$
We write $K^{d}$ (resp., $K^{f}$) to denote the strike price denominated in the domestic currency (resp., the foreign currency). It is assumed that call and put options with payoffs $C_{T}(S^{d},K^{d},T)=(S^{d}_{T}-K^{d})^{+}$ and $P_{T}(S^{d},K^{d},T)=(K^{d}-S^{d}_{T})^{+}$ for all strikes $K^{d}$ are traded, which is a natural assumption given that the portfolio $S^{d}$ typically represents a market index. In the case of a bespoke portfolio $S^{d}$, one may use the Black-Scholes formula to obtain model prices of calls and puts.
The case of a domestic EPS does not require a detailed study since it was already examined in Xu et al. [19]. Let us only give the following result, which is an easy consequence of Proposition 2.1.
Corollary 3.1.
The static hedging in terms of domestic equity options for the domestic EPS is given by Proposition 2.1 with $\psi=\psi_{d},N=N^{d}$ and $X=S^{d}$, that is, $\mbox{\rm EPS\,}^{d}_{T}=\widehat{H}^{d}_{T}$ where for every $t\in[0,T]$,
$\widehat{H}^{d}_{t}=N^{d}\sum_{i=0}^{n}\frac{p_{i+1}-p_{i}}{S^{d}_{0}}\,P_{t}(S^{d},K^{l}_{i}(S^{d}_{0}),T)-N^{d}\sum_{j=0}^{m}\frac{f_{j+1}-f_{j}}{S^{d}_{0}}\,C_{t}(S^{d},K^{g}_{j}(S^{d}_{0}),T)$
where $K^{l}_{i}(S^{d}_{0})=(1+l_{i})S^{d}_{0}$ and $K^{g}_{j}(S^{d}_{0})=(1+g_{j})S^{d}_{0}$ for every $i=0,1,\dots,n$ and $j=0,1,\dots,m$.
If we assume that the dynamics of the domestic portfolio $S^{d}$ is given by (3.2), then we may use the classical Black-Scholes options pricing formula to obtain the following result for domestic EPSs. It is worth stressing that plain-vanilla options on typical domestic and foreign portfolios, $S^{d}$ and $S^{f}$, are likely to be actively traded on options exchanges and thus Corollary 3.1, which is a model-free result, is sufficient for real-life pricing and hedging of a domestic EPS. Corollary 3.2 is useful when studying model pricing of domestic EPSs and using their model prices in the context of more complex cross-currency EPSs, which are studied in foregoing sections. We henceforth write $\Phi$ to denote the cumulative density function of the standard normal distribution.
Corollary 3.2.
The price at time 0 of the domestic EPS with the payoff $\mbox{\rm EPS\,}^{d}_{T}$ at time $T$ equals
$\mbox{\rm EPS\,}^{d}_{0}$ $=N^{d}\sum_{i=0}^{n}(p_{i+1}-p_{i})\Big{(}e^{-r^{d}T}(1+l_{i})\Phi\big{(}-d_{-}(l_{i},T)\big{)}-\Phi\big{(}-d_{+}(l_{i},T)\big{)}\Big{)}$
$-N^{d}\sum_{j=0}^{m}(f_{j+1}-f_{j})\Big{(}\Phi\big{(}d_{+}(g_{j},T)\big{)}-e^{-r^{d}T}(1+g_{j})\Phi\big{(}d_{-}(g_{j},T)\big{)}\Big{)}$
where
$d_{\pm}(x,T):=\frac{1}{\|\bm{\sigma}^{d}\|\sqrt{T}}\bigg{[}-\ln(1+x)+\Big{(}r^{d}\pm\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}\Big{)}T\bigg{]}.$
Analogous results can be formulated for the nominal foreign EPS with the payoff denominated in the foreign currency and given by
$\mbox{\rm EPS\,}^{f}_{T}:=N^{f}\psi_{f}(R^{f}_{T})=(Q_{0})^{-1}(1-w)N\psi_{f}(R^{f}_{T})$
so that
$\mbox{\rm EPS\,}^{f}_{T}=N^{f}\sum_{i=0}^{n}\frac{p_{i+1}-p_{i}}{S^{f}_{0}}\,P_{T}(S^{f},K^{l}_{i}(S^{f}_{0}),T)-N^{f}\sum_{j=0}^{m}\frac{f_{j+1}-f_{j}}{S^{f}_{0}}\,C_{T}(S^{f},K^{g}_{j}(S^{f}_{0}),T).$
Similarly, to obtain a foreign market counterpart of Corollary 3.2, it suffices to substitute $S^{d}$ and $\bm{\sigma}^{d}$ with $S^{f}$ and $\bm{\sigma}^{f}$, respectively, to obtain the price $\mbox{\rm EPS\,}^{f}_{0}$ in the foreign currency. Then the domestic payoff of the nominal foreign EPS equals $Q_{T}N^{f}\psi_{f}(R^{f}_{T})$ and its price in the domestic currency is obtained by simply multiplying its foreign price $\mbox{\rm EPS\,}^{f}_{0}$ by the exchange rate $Q_{0}$. Needless to say, all parameters of domestic and foreign EPSs are completely arbitrary, for instance, a domestic buffer EPS can be combined with a foreign floor EPS with different protection and fee rates.
Remark 3.1.
Typically, the domestic and foreign portfolios of investors can be linked to existing market indices, such as the S&P 500 Index in the U.S. or the S&P/ASX 200 Index in Australia, which both cover a large majority of listed equities in respective economies. It is worth noting that many financial institutions maintain and offer shares in funds containing multiple stocks which are designed to meet some goal. These funds are referred to as exchange traded funds (ETFs) and, most frequently, they are designed to approximate the performance of market indices. Therefore, they can be considered as actively traded proxies of market indices and thus the assumption that market indices can be seen as traded assets is innocuous.
Furthermore, it is clear from Corollary 3.1 that liquidly traded options on market indices, which are available on major exchanges worldwide, can be used by providers of domestic and foreign EPSs for implementation of static hedging and hence also pricing of various EPS products they offer, not only at the time of their issuance but also for mark-to-market purposes if cancellation of an EPS by its holder is allowed. In this context, it is worth mentioning the CBOE S&P 500 Index options (SPX), Nanos S&P 500 Index options (NANOS), and SPDR S&P 500 ETF Options (SPY) offered in the U.S., as well as ASX S&P/ASX 200 Index options (XJO) actively traded in Australia. Note that the abovementioned index options are of European style, except for SPY options, which are of American style.
3.4 Protection Against Effective Foreign Losses
If the holder aims to get protection against losses from the foreign part of his portfolio with the initial wealth $N$ and the weight $1-w$ for foreign assets, he may enter into the effective foreign EPS with notional principal $N^{f,d}$. Then the payoff at time $T$ to the provider becomes
$\mbox{\rm EPS\,}^{f,e}_{T}:=N^{f,d}\psi_{f}(R^{f,e}_{T})=N^{f}Q_{0}\psi_{f}(R^{f,e}_{T})=(1-w)N\psi_{f}(R^{f,e}_{T}).$
In order to distinguish the different structures of these two different EPS products represented by $\psi_{d}$ and $\psi_{f}$, one may use the notation $l_{i},g_{j},p_{i},f_{j}$ for the domestic EPS and $l^{f}_{i},g^{f}_{j},p^{f}_{i},f^{f}_{j}$ for the foreign EPS but we believe that it would be to heavy and hence the domestic notation is used for both cases.
For the purpose of static hedging of an effective foreign EPS, we assume that the provider can trade in foreign equity options with the foreign equity struck in domestic currency as the underlying asset and strike price given in the domestic currency. Notice that the price of such options is given in the domestic currency since they are offered to domestic investors by Australian finance companies on OTC market. By definition, a holder of a foreign equity call option struck in the domestic currency has the right to buy at time $T$ one unit of $S^{f}$ by paying $K$ units of domestic currency, which formally gives the payoff equal to $(S^{f,e}_{T}-K^{d})^{+}$ in the domestic currency. We denote $\widetilde{P}_{T}(S^{f,e},K^{d},T)=(K^{d}-S^{f,e}_{T})^{+}$ and $\widetilde{C}_{T}(S^{f,e},K^{d},T)=(S^{f,e}_{T}-K^{d})^{+}$. Similarly, let $\widetilde{P}(S^{f,e},K^{d},T)$ and $\widetilde{C}(S^{f,e},K^{d},T)$ be the prices foreign equity put and call options on the synthetic risky asset $S^{f,e}$ represented the foreign reference portfolio struck in the domestic currency.
Corollary 3.3.
The static hedging in terms of foreign equity call and put options struck in the domestic currency for an effective foreign EPS is given by Proposition 2.1 with $\psi=\psi_{f},N=N^{f,d}$ and $X=S^{f,e}$, that is, $\mbox{\rm EPS\,}^{f,e}_{T}=\widehat{H}^{f,e}_{T}$ where, for every $t\in[0,T]$,
$\widehat{H}^{f,e}_{t}=N^{f,d}\sum_{i=0}^{n}\frac{p_{i+1}-p_{i}}{S^{f,e}_{0}}\,\widetilde{P}_{t}(S^{f,e},K^{l}_{i}(S^{f,e}_{0}),T)-N^{f,d}\sum_{j=0}^{m}\frac{f_{j+1}-f_{j}}{S^{f,e}_{0}}\,\widetilde{C}_{t}(S^{f,e},K^{g}_{j}(S^{f,e}_{0}),T)$
where $K^{l}_{i}(S^{f,e}_{0})=(1+l_{i})S^{f,e}_{0}$ and $K^{g}_{j}(S^{f,e}_{0})=(1+g_{j})S^{f,e}_{0}$ for every $i=0,1,\dots,n$ and $j=0,1,\dots,m$.
Remark 3.2.
Several ETFs available in Australia allow to track the performance of the S&P 500 AUD Index, which is designed to measure the AUD performance of large capitalisation U.S. equities. Therefore, investing in such ETFs can be seen as a proxy for the theoretical concept of the value process $S^{f,e}$ representing the domestic price of a foreign portfolio, e.g., a particular market index in the foreign economy. Corollary 3.3 shows that related options of European style could be used to hedge and price effective foreign EPSs and thus, in principle, their current market values would be given in terms of market quotes for options on such ETFs provided that they are publicly available.
It is a realistic assumption that the above mentioned options are available on the OTC market, in which case the provider of an effective foreign EPS can establish the static hedging strategy. However, for the purpose of getting a better insight into properties of arbitrage-free prices of effective foreign EPSs of various kinds and with differing parameters, we will also use the following classical result, which is known to hold within the setup of Section 3.1. In fact, it is an easy consequence of the dynamics of $S^{f,e}$ under ${{\mathbb{Q}}}^{d}$, which read
$dS^{f,e}_{t}=S^{f,e}_{t}\big{(}r^{d}\,dt+(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\textbf{W}_{t}^{d}\big{)}.$
Proposition 3.1.
The price a foreign call option struck in the domestic currency with the terminal payoff at maturity $T$ equal to $\widetilde{C}_{T}(S^{f,e},K^{d},T)=(S^{f,e}_{T}-K^{d})^{+}$ is given by
$\widetilde{C}_{t}(S^{f,e},K^{d},T)=S^{f,e}_{t}\Phi\big{(}\widetilde{d}_{+}(S^{f,e}_{t},T-t)\big{)}-\widetilde{K}^{d}_{t}\Phi\big{(}\widetilde{d}_{-}(S^{f,e}_{t},T-t)\big{)}$
where $\widetilde{K}^{d}_{t}:=e^{-r^{d}(T-t)}K^{d}$ and
$\widetilde{d}_{\pm}(S^{f,e}_{t},T-t):=\frac{1}{\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|\sqrt{T-t}}\bigg{[}\ln\frac{S^{f,e}_{t}}{\widetilde{K}^{d}_{t}}\pm\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2}(T-t)\bigg{]}.$
The price of a put option with the payoff $\widetilde{P}_{T}(S^{f,e},K^{d},T)=(K^{d}-S^{f,e}_{T})^{+}$ equals
$\widetilde{P}_{t}(S^{f,e},K^{d},T)=\widetilde{K}^{d}_{t}\Phi\big{(}-\widetilde{d}_{-}(S^{f,e}_{t},T-t)\big{)}-S^{f,e}_{t}\Phi\big{(}-\widetilde{d}_{+}(S^{f,e}_{t},T-t)\big{)}.$
Notice that Proposition 3.1 allows for dynamic hedging of an effective foreign EPS using the underlying portfolio $S^{f}$ (formally represented by $S^{f,e}$) and the domestic money market account $B^{d}$. Furthermore, by combining Corollary 3.3 with Proposition 3.1, we obtain the pricing formula for an effective foreign EPS, which holds within the setup of Section 3.2.
Corollary 3.4.
The price at time 0 of the effective foreign EPS with the payoff $\mbox{\rm EPS\,}^{f}_{T}$ at time $T$ equals
$\mbox{\rm EPS\,}^{f,e}_{0}$ $=N^{f,d}\sum_{i=0}^{n}(p_{i+1}-p_{i})\Big{(}e^{-r^{d}T}(1+l_{i})\Phi\big{(}-k_{-}(l_{i},T)\big{)}-\Phi\big{(}-k_{+}(l_{i},T)\big{)}\Big{)}$
$-N^{f,d}\sum_{j=0}^{m}(f_{j+1}-f_{j})\Big{(}\Phi\big{(}k_{+}(g_{j},T)\big{)}-e^{-r^{d}T}(1+g_{j})\Phi\big{(}k_{-}(g_{j},T)\big{)}\Big{)}$
where
$k_{\pm}(x,T):=\frac{1}{\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|\sqrt{T}}\bigg{[}-\ln(1+x)+\Big{(}r^{d}\pm\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2}\Big{)}T\bigg{]}.$
3.5 Protection Against Quanto Foreign Losses
A quantity-adjusting derivative (or, simply, a quanto derivatinve) is a cross-currency derivative in which the underlying asset is denominated in one currency but settlement is made in another currency at a predetermined exchange rate, hereafter denoted as $\overline{Q}$. Another name for such a financial product is a guaranteed exchange rate derivative since they enable investors to fix at the contract’s inception the exchange rate between two currencies. By construction, quanto derivatives help to safeguard the investors against unfavourable fluctuations in the exchange rate but, of course, at the cost of no benefiting from a favourable move of the exchange rate.
In the quanto EPS with the predetermined exchange rate $\overline{Q}$, the provider’s cash flow at time $T$ equals, in the domestic currency,
$\mbox{\rm EPS\,}^{f,q}_{T}:=\overline{Q}N^{f}\psi_{f}(R^{f}_{T}).$
Let the quanto call and put options written on $S^{f}$ be denoted by $\overline{C}_{T}(S^{f},K^{f},T,\overline{Q})=\overline{Q}(S^{f}_{T}-K^{f})^{+}$ and $\overline{P}_{T}(S^{f},K^{f},T,\overline{Q})=\overline{Q}(K^{f}-S^{f}_{T})^{+}$ where $S^{f}_{T}$ and $K^{f}$ are given in the foreign currency but the option’s payoff is expressed in the domestic currency. Similarly, $\overline{C}(S^{f},K^{f},T,\overline{Q})$ and $\overline{P}(S^{f},K^{f},T,\overline{Q})$ stand for the price processes of quanto call and put options written on $S^{f}$. We first state a model-free result on static hedging of a quanto EPS and hence also the price in terms of quanto options.
Corollary 3.5.
The static hedging in terms of quanto options for a quanto EPS is given by Proposition 2.1 with $\psi=\psi_{f},N=N^{f}$ and $X=S^{f}$, that is, $\mbox{\rm EPS\,}^{f,q}_{T}=\widehat{H}^{f,q}_{T}$ where, for every $t\in[0,T]$,
$\widehat{H}^{f,q}_{t}=N^{f}\sum_{i=0}^{n}\frac{p_{i+1}-p_{i}}{S^{f}_{0}}\,\overline{P}_{t}(S^{f},K^{l}_{i}(S^{f}_{0}),T,\overline{Q})-N^{f}\sum_{j=0}^{m}\frac{f_{j+1}-f_{j}}{S^{f}_{0}}\,\overline{C}_{t}(S^{f},K^{g}_{j}(S^{f}_{0}),T,\overline{Q})$
where $K^{l}_{i}(S^{f}_{0})=(1+l_{i})S^{f}_{0}$ and $K^{g}_{j}(S^{f}_{0})=(1+g_{j})S^{f}_{0}$ for every $i=0,1,\dots,n$ and $j=0,1,\dots,m$.
Remark 3.3.
Similar to Remark 3.2, as a proxy for the concept of the quanto foreign return, we could mention returns on exchange traded funds aiming to provide investors with the performance of the S&P 500 AUD Hedged Index, which uses a currency hedged strategy to reduce the risk of currency fluctuations while accepting underlying market risk of major U.S. equities. As usual, we refer to AUD as the domestic currency but analogous indices are also offered for other major currencies. Then, as shown in Corollary 3.5, the market value of a quanto EPS would be given in terms of market prices of options on such ETFs, of course, provided that they are traded on the OTC market.
If the provider cannot directly trade in quanto options, then a plausible alternative is to use a stochastic model to compute their prices and obtain a synthetic versions of quanto options by dynamic trading. To this end, one may use the following result, which holds in the model of Section 3.1. Let $\overline{Q}$ be a predetermined exchange rate and let us denote $\delta^{q}:=r^{f}-r^{d}-\bm{\sigma}^{f}(\bm{\sigma^{q}})^{*}$.
Proposition 3.2.
The price of a quanto call option with the payoff $\overline{C}_{T}(S^{f},K^{f},T,\overline{Q})=\overline{Q}(S^{f}_{T}-K^{f})^{+}$ in the domestic currency is given by
$\overline{C}_{t}(S^{f},K^{f},T,\overline{Q})=\overline{Q}\Big{(}S^{f}_{t}e^{\delta^{q}(T-t)}\Phi\big{(}\bar{d}_{+}(S^{f}_{t},T-t)\big{)}-\widetilde{K}^{f}_{t}\Phi\big{(}\bar{d}_{-}(S^{f}_{t},T-t)\big{)}\Big{)}$
where $\widetilde{K}^{f}_{t}=e^{-r^{d}(T-t)}K^{f}$ and
$\bar{d}_{\pm}(S^{f}_{t},T-t)=\frac{1}{\|\bm{\sigma}^{f}\|\sqrt{T-t}}\bigg{[}\ln\frac{S^{f}_{t}}{\widetilde{K}^{f}_{t}}+\Big{(}\delta^{q}\pm\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}\Big{)}(T-t)\bigg{]}.$
The price of a quanto put option with the payoff $\overline{P}_{T}(S^{f},K^{f},T,\overline{Q})=\overline{Q}(K^{f}-S^{f}_{T})^{+}$ equals
$\overline{P}_{t}(S^{f},K^{f},T,\overline{Q})=\overline{Q}\Big{(}\widetilde{K}^{f}_{t}\Phi\big{(}-\bar{d}_{-}(S^{f}_{t},T-t)\big{)}-S^{f}_{t}e^{\delta^{q}(T-t)}\Phi\big{(}-\bar{d}_{+}(S^{f}_{t},T-t)\big{)}\Big{)}.$
From Corollary 3.5 and Proposition 3.2, we obtain the pricing formula for a quanto EPS in a model of Section 3.2. Notice that we now set $\overline{Q}=Q_{0}$ and thus the equality $\overline{Q}N^{f}=N^{f,d}$ holds.
Corollary 3.6.
The price at time 0 of a quanto foreign EPS with the payoff $\mbox{\rm EPS\,}^{f}_{T}$ at time $T$ with $\overline{Q}=Q_{0}$ equals
$\mbox{\rm EPS\,}^{f,q}_{0}$ $=N^{f,d}\sum_{i=0}^{n}(p_{i+1}-p_{i})\Big{(}e^{-r^{d}T}(1+l_{i})\Phi\big{(}-\bar{k}_{-}(l_{i},T)\big{)}-e^{\delta^{q}T}\Phi\big{(}-\bar{k}_{+}(l_{i},T)\big{)}\Big{)}$
$-N^{f,d}\sum_{j=0}^{m}(f_{j+1}-f_{j})\Big{(}e^{\delta^{q}T}\Phi\big{(}\bar{k}_{+}(g_{j},T)\big{)}-e^{-r^{d}T}(1+g_{j})\Phi\big{(}\bar{k}_{-}(g_{j},T)\big{)}\Big{)}$
where
$\bar{k}_{\pm}(x,T)=\frac{1}{\|\bm{\sigma}^{f}\|\sqrt{T}}\bigg{[}-\ln(1+x)+\Big{(}r^{f}-\bm{\sigma}^{f}(\bm{\sigma^{q}})^{*}\pm\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}\Big{)}T\bigg{]}.$
4 Pricing and Hedging of an Aggregated Effective EPS
In Section 3, it was assumed that the EPS provider can directly participate in the foreign market to fully hedge the nominal return on the foreign reference portfolio. However, we do not expect that this situation will be typical in the real world since it is more likely that holders of a superannuation account will pay more attention to the effective return on their foreign holdings expressed in the domestic currency and pooled with their domestic holdings. The holder can choose to invest in global markets and can track the return on the foreign portfolio but his account uses exclusively the domestic currency in the final settlement and thus the provider should also offer EPSs on the foreign equity return struck in the domestic currency in order to satisfy the holder’s demand for protection.
Therefore, after discussing separate hedging strategies for domestic equity return and foreign equity return, we now wish to examine the joint hedging strategy for a total return on the whole reference portfolio. For instance, when considering the pension system, the rate of return on the pension account is obtained by merging returns on various assets in the pension portfolio together and calculating the total return. Although the pension holder knows the proportion of each asset in the whole pension portfolio, it is hard to separate the returns on each asset from the total return. Obviously, focusing on the total return on the cross-currency reference portfolio and hedging the total return directly would be a better choice for both the EPS buyer and provider. In this situation, the insurance company’s consideration and hedging strategy building method is different from a separate hedging strategy for domestic and foreign equity, and it is more difficult to build a joint hedging portfolio than a separate hedging strategy.
We will now examine the case of a cross-currency effective EPS on aggregated returns. As usual, $N$ is the nominal value in the domestic currency of the holder’s portfolio of domestic and foreign assets and also the notional principal of an EPS under study.
We first consider the effective wealth of a cross-currency portfolio, that is, its total wealth expressed in the domestic currency. For every notional principal $N>0$ and weight $w\in[0,1]$, the aggregated effective wealth $X^{w,e}$ of a cross-currency portfolio satisfies $X^{w,e}_{0}=N=\alpha_{0}S^{d}_{0}+\beta_{0}S^{f,e}_{0}$ and, for every $t\in[0,T]$,
$X^{w,e}_{t}=\alpha_{0}S^{d}_{t}+\beta_{0}S^{f,e}_{t}=N(wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f,e}_{t}(S^{f,e}_{0})^{-1})=NS^{w,e}_{t}$
where we denote
$S^{w,e}_{t}:=N^{-1}X^{w,e}_{t}=wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f,e}_{t}(S^{f,e}_{0})^{-1}=w(1+R^{d}_{t})+(1-w)(1+R^{f,e}_{t})$
so that, in particular, $S^{w,e}_{0}=1$. We now consider the process $S^{w,e}$ as a reference portfolio for an EPS and thus also as the underlying process for related cross-currency call and put options. Of course, the process $S^{w,e}$ does not necessarily represent a listed stock or a quoted market index, although in some situations options on $S^{w,e}$ (or, equivalently, on its returns) could be available on OTC market. Therefore, as opposed to the previous section, we will now focus on pricing and hedging of EPSs on aggregated wealth within the framework of the cross-currency model. For any notional principal $N>0$ and weight $w\in[0,1]$, the aggregated effective rate of return $R^{w,e}$ on a cross-currency portfolio equals, and $t\in[0,T]$,
$R^{w,e}_{t}:=\frac{X^{w,e}_{t}-X^{w,e}_{0}}{X^{w,e}_{0}}=S^{w,e}_{t}-1=wR^{d}_{t}+(1-w)R^{f,e}_{t}.$
Definition 4.1.
In the cross-currency effective EPS, the provider’s cash flow at time $T$ is given in the domestic currency and equals
$\mbox{\rm EPS\,}^{w,e}_{T}:=N\psi_{a}(R^{w,e}_{T})=N\psi_{a}\big{(}wR^{d}_{T}+(1-w)R^{f,e}_{T}\big{)}.$
To derive the static hedging strategy for the cross-currency effective EPS, we consider cross-currency basket call and put options written on $S^{w,e}$ with the payoffs $\widehat{C}_{T}(S^{w,e},K^{d},T):=(S^{w,e}_{T}-K^{d})^{+}$ and $\widehat{P}_{T}(S^{w,e},K^{d},T):=(K^{d}-S^{w,e}_{T})^{+}$ where the payoffs are given in the domestic currency.
Corollary 4.1.
The static hedging in terms of options written on $S^{w,e}$ for the aggregated effective EPS is given by Proposition 2.1 with $\psi=\psi_{a}$ and $X=S^{w,e}$, that is, $\mbox{\rm EPS\,}^{w,e}_{T}=\widehat{H}^{w,e}_{T}$ where, for every $t\in[0,T]$,
$\widehat{H}^{w,e}_{t}=N\sum_{i=0}^{n}(p_{i+1}-p_{i})\widehat{P}_{t}(S^{w,e},k^{l}_{i},T)-N\sum_{j=0}^{m}(f_{j+1}-f_{j})\widehat{C}_{t}(S^{w,e},k^{g}_{j},T)$
where $k^{l}_{i}=1+l_{i}$ and $k^{g}_{j}=1+g_{j}$ for every $i=0,1,\dots,n$ and $j=0,1,\dots,m$.
Using this static hedging strategy, the provider can perfectly hedge the effective return on the whole reference portfolio, which includes both domestic equities and foreign equities struck in the domestic currency. Although the provider’s hedging strategy is fairly simple, it hinges on a strong postulate that the relevant call and put options on the value process $S^{w,e}$ are actively traded, which is a rather unrealistic postulate, in particular, due to the fact that $S^{w,e}$ depends on the choice of the weight $w$. Therefore, we will also examine the theoretical pricing and hedging of an aggregated effective EPS within the framework of the market model introduced in Section 3.2, as well as its superhedging using separate domestic and foreign EPSs.
To this end, we start by noticing that
$S^{w,e}_{t}=wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f,e}_{t}(S^{f,e}_{0})^{-1}=w\widehat{S}^{d}_{t}+(1-w)\widehat{S}^{f,e}_{t}$
where, for the sake of computational convenience, we introduce normalized processes $\widehat{S}^{d}_{t}:=S^{d}_{t}(S^{d}_{0})^{-1}$ and $\widehat{S}^{f,e}_{t}:=S^{f,e}_{t}(S^{f,e}_{0})^{-1}$ so that $\widehat{S}^{d}_{0}=\widehat{S}^{f,e}_{0}=1$. Of course, that normalization does not affect the dynamics of price processes $S^{d}$ and $S^{f,e}$, as specified in Section 3.2, but only their initial values and thus we have that
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}r^{d}\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}\big{)},$
$d\widehat{S}^{f,e}_{t}=\widehat{S}^{f,e}_{t}\big{(}r^{d}\,dt+(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\textbf{W}_{t}^{d}\big{)}.$
The cross-currency basket call option written on $S^{w,e}$ has the terminal payoff at maturity $T$ given by $C_{T}(S^{w,e},K^{d},T):=(S^{w,e}_{T}-K^{d})^{+}$ in the domestic currency. Since $S^{w,e}_{T}$ is a sum of lognormally distributed random variables it does not follow a lognormal distribution and thus the Black-Scholes-like pricing formula for a basket option is not available. However, there are several other ways of computing, at least approximately, the prices of basket options:
(i) approximate analytical pricing using the geometric averaging method,
(ii) approximate analytical pricing using the moment matching technique,
(iii) numerical pricing using Monte-Carlo simulation.
4.1 Approximate Pricing using Geometric Averaging Method
For the analytical approximation of the price of the cross-currency basket call option, it is convenient to rewrite the pricing equality as follows
$\widehat{C}_{0}(S^{w,e},K^{d},T)=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}(S^{w,e}_{T}-K^{d})^{+}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}w\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f,e}_{T}-\widetilde{K}^{d}\big{)}^{+}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{w,e}_{T}-\widetilde{K}^{d}\big{)}^{+}\big{]}$
where we introduce the discounted prices $\widetilde{S}^{d}_{t}:=e^{-r^{d}t}\widehat{S}^{d}_{t}$ and $\widetilde{S}^{f,e}_{t}:=e^{-r^{d}t}\widehat{S}^{f,e}_{t}$ and $\widetilde{K}^{d}:=e^{-r^{d}T}K^{d}$ so that
$d\widetilde{S}^{d}_{t}=\widetilde{S}^{d}_{t}\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d},$
$d\widetilde{S}^{f,e}_{t}=\widetilde{S}^{f,e}_{t}(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\textbf{W}_{t}^{d},$
with $\widetilde{S}^{d}_{0}=\widetilde{S}^{f,e}_{0}=1$. To get an explicit approximation for the pricing formula of a classical basket option on equities, it was proposed by Gentle [8] to use an approximation of the weighted arithmetic mean $\widetilde{S}^{w,e}_{T}:=w\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f,e}_{T}$ by a similarly weighted geometric mean, which is henceforth denoted as $G_{T}$, and given by
$G_{T}:=\big{(}\widetilde{S}^{d}_{T}\big{)}^{w}\big{(}\widetilde{S}^{f,e}_{T}\big{)}^{(1-w)}=e^{(w\bm{\sigma}^{d}+(1-w)(\bm{\sigma}^{f}+\bm{\sigma}^{q}))\textbf{W}_{T}^{d}-\frac{1}{2}\gamma T}=e^{\xi_{T}-\frac{1}{2}\gamma T}$
where
$\gamma:=w\|\sigma^{d}\|^{2}+(1-w)\|\sigma^{f}+\sigma^{q}\|^{2}$
and $\xi_{T}$ is a Gaussian random variable under ${{\mathbb{Q}}}^{d}$ with null expected value and the variance $v^{2}T$ where
$v^{2}:=\|w\bm{\sigma}^{d}+(1-w)(\bm{\sigma}^{f}+\bm{\sigma}^{q})\|^{2}=w^{2}\|\sigma^{d}\|^{2}+2w(1-w)\sigma^{d}(\sigma^{f}+\sigma^{q})^{*}+(1-w)^{2}\|\sigma^{f}+\sigma^{q}\|^{2}.$
It is clear that the random variable $G_{T}$ has the lognormal distribution under ${{\mathbb{Q}}}^{d}$ with the expected value
${{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}(G_{T})$ $=e^{\frac{1}{2}(v^{2}-\gamma)T}=\exp\Big{(}-\frac{1}{2}w(1-w)\big{(}\|\sigma^{d}\|^{2}-2\sigma^{d}(\sigma^{f}+\sigma^{q})^{*}+\|\sigma^{f}+\sigma^{q}\|^{2}\big{)}T\Big{)}$
$=\exp\Big{(}-\frac{1}{2}w(1-w)\|\sigma^{d}-\sigma^{f}-\sigma^{q}\|^{2}T\Big{)}=:\lambda.$
Then the proposed approximated value for the price of the cross-currency basket call option is given by the expected value
$C^{G}_{0}(\widehat{K}^{d},T):={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}G_{T}-\widehat{K}^{d}\big{)}^{+}\big{]}$
where $\widehat{K}^{d}$ equals
$\widehat{K}^{d}:=\widetilde{K}^{d}+{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}G_{T}-\big{(}w_{t}\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f,e}_{T}\big{)}\big{]}=\widetilde{K}^{d}+\lambda-1$
where in turn we used the equality
${{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}w\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f,e}_{T}\big{]}=w\widetilde{S}^{d}_{0}+(1-w)\widetilde{S}^{f,e}_{0}=1,$
which is a consequence of the martingale property of $\widetilde{S}^{d}$ and $\widetilde{S}^{f,e}$ under ${{\mathbb{Q}}}^{d}$. Elementary computations now lead to the following result, which is a variant of a more general pricing formula for an option written on a basket on $n$ domestic equities (see Lemma 6.9.1 in Section 6.9 of Musiela and Rutkowski [16]).
Proposition 4.1.
The price $\widehat{C}_{0}(S^{w,e},K^{d},T)$ of the cross-currency basket call option can be approximated by
$C^{G}_{0}(\widehat{K}^{d},T)=\lambda\Phi\big{(}\widehat{d}_{+}(\widehat{K}^{d},T)\big{)}-\widehat{K}^{d}\Phi\big{(}\widehat{d}_{-}(\widehat{K}^{d},T)\big{)}$
where
$\widehat{d}_{\pm}(\widehat{K}^{d},T)=\frac{1}{v\sqrt{T}}\Big{[}\ln\frac{\lambda}{\widehat{K}^{d}}\pm\frac{1}{2}v^{2}T\Big{]}.$
The price $\widehat{P}_{0}(S^{w,e},K,T)$ of the cross-currency basket put option with the payoff $\widehat{P}_{T}(S^{w,e},K^{d},T)=(K^{d}-S^{w,e}_{T})^{+}$ can be approximated by
$P^{G}_{0}(\widehat{K}^{d},T)=\widehat{K}^{d}\Phi\big{(}-\widehat{d}_{-}(\widehat{K}^{d},T)\big{)}-\lambda\Phi\big{(}-\widehat{d}_{+}(\widehat{K}^{d},T)\big{)}.$
Remark 4.1.
The put-call parity for cross-currency basket options has the form $\widehat{C}_{0}(S^{w,e},K^{d},T)-\widehat{P}_{0}(S^{w,e},K^{d},T)=1-\widetilde{K}^{d}$ (recall that $\widetilde{S}^{w,e}$ is a martingale under ${\mathbb{Q}}^{d}$ and $\widetilde{S}^{w,e}_{0}=1$) whereas for the approximated values it is given by $C^{G}_{0}(\widehat{K}^{d},T)-P^{G}_{0}(\widehat{K}^{d},T)=\lambda-\widehat{K}^{d}=1-\widetilde{K}^{d}$. Therefore, as was expected, it suffices to focus on the pricing formula for a cross-currency basket call option.
4.2 Approximate Pricing using Moment Matching Method
Borovkova et al. [3] consider basket option pricing problems in futures markets. They use shifted lognormal models with three-moment matching method to approximate baskets with possibly negative weights, resulting in fairly accurate approximation for prices of basket options. To be more specific, the method uses the first three moments to match the four real-valued parameters $m,s,\tau,c$ of the approximating random variable
$c(e^{sZ+m}+\tau)$
where $Z$ is a standard normal random variable under a probability measure ${\mathbb{Q}}$, $e^{m}$ can be interpreted as a scale parameter, $s$ is a shape parameter, and $\tau$ is a location parameter. The parameter $c=\pm 1$ is included to adjust the overall sign of the basket price when weights are not all of the same sign and it is chosen to coincide with the sign of the skewness of the basket, which is henceforth denoted as $\eta_{T}$ (see equation (6)).
With the algorithm developed in [3], a real-valued closed-form solution for the set of nonlinear equations is given by Hu et al. [10], and the approximate closed-form basket option price is expressed as an explicit functional form of the basket’s mean, variance, and skewness. We start by computing the first three moments of the basket based on stocks. In order to ensure the comprehensiveness and independence of our paper, we have included these computations in the following Lemma that starts with changing the labels to a simpler form,
Recall from Section 4.1 that we wish to compute
$\widehat{C}_{0}(S^{w,e},K^{d},T)={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}w\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f,e}_{T}-\widetilde{K}^{d}\big{)}^{+}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{w,e}_{T}-\widetilde{K}^{d}\big{)}^{+}\big{]}$
where $\widetilde{K}^{d}:=e^{-r^{d}T}K^{d}$ and $\widetilde{S}^{w,e}_{t}:=w\widetilde{S}^{d}_{t}+(1-w)\widetilde{S}^{f,e}_{t}$. Since
$d\widetilde{S}^{d}_{t}=\widetilde{S}^{d}_{t}\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d},$
$d\widetilde{S}^{f,e}_{t}=\widetilde{S}^{f,e}_{t}(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\textbf{W}_{t}^{d},$
with $\widetilde{S}^{d}_{0}=\widetilde{S}^{d}_{0}=1$ we have that
$\widetilde{S}^{w,e}_{T}=we^{\bm{\sigma}^{d}\textbf{W}_{T}^{d}-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}+(1-w)e^{(\bm{\sigma}^{f}+\bm{\sigma}^{q})\sqrt{T}\textbf{W}_{T}^{d}-\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2})T}.$
Lemma 4.1.
Let $w_{1}=w,\,w_{2}=1-w,\,\bm{\sigma}_{1}=\bm{\sigma}^{d}$ and $\bm{\sigma}_{2}=\bm{\sigma}^{f}+\bm{\sigma}^{q}$. Then the first three moments of the random variable $\widetilde{S}^{w,e}_{T}$ under the martingale measure ${\mathbb{Q}}^{d}$ are
$\widetilde{M}_{1}$ $=w_{1}+w_{2}=1,$
$\widetilde{M}_{2}$ $=w_{1}^{2}e^{\|\bm{\sigma}_{1}\|^{2}T}+2w_{1}w_{2}e^{\bm{\sigma}_{1}\bm{\sigma}_{2}^{*}T}+w_{2}^{2}e^{\|\bm{\sigma}_{2}\|^{2}T},$
$\widetilde{M}_{2}$ $=w_{1}^{3}e^{3\|\bm{\sigma}_{1}\|^{2}T}+3w_{1}^{2}w_{2}e^{\big{(}\|\bm{\sigma}_{1}\|^{2}+2\bm{\sigma}_{1}\bm{\sigma}_{2}^{*}\big{)}T}+3w_{1}w_{2}^{2}e^{\big{(}\|\bm{\sigma}_{2}\|^{2}+2\bm{\sigma}_{1}\bm{\sigma}_{2}^{*}\big{)}T}+w_{2}^{3}e^{3\|\bm{\sigma}_{2}\|^{2}T}.$
We denote the mean of $\widetilde{S}^{w,e}_{T}$ by $\mu_{T}:=\widetilde{M}_{1}=1$, its standard deviation by $\sigma_{T}^{2}:=\widetilde{M}_{2}-\mu_{T}^{2}$, and its skewness by
$\eta_{T}:=\frac{\widetilde{M}_{3}-3\mu_{T}\sigma_{T}^{2}-\mu_{T}^{3}}{\sigma_{T}^{3}}.$ (6)
The first three moments of the random variable $c(e^{sZ+m}+\tau)$ are
$M_{1}(c,s,m,\tau)$ $:=c\big{(}e^{\frac{1}{2}s^{2}+m}+\tau\big{)},$
$M_{2}(c,s,m,\tau)$ $:=c^{2}\big{(}e^{2s^{2}+2m}+2\tau e^{\frac{1}{2}s^{2}+m}+\tau^{2}\big{)},$ (7)
$M_{3}(c,s,m,\tau)$ $:=c^{3}\big{(}e^{\frac{9}{2}s^{2}+3m}+3\tau e^{2s^{2}+2m}+3\tau^{2}e^{\frac{1}{2}s^{2}+m}+\tau^{3}\big{)}.$
where the parameters $c,s,m,\tau,$ can be obtained by solving the moment matching conditions $\widetilde{M}_{i}=M_{i}(c,s,m,\tau)$ for $i=1,2,3$. According to Hu et al. [10], we have the following result.
Lemma 4.2.
One set of real-valued solution $c,s,m,\tau$ that satisfy the moment matching conditions $\widetilde{M}_{i}=M_{i}(c,s,m,\tau)$ for $i=1,2,3$ is given by
$c=\text{sgn}(\eta_{T}),\quad s=(\ln(x))^{\frac{1}{2}},\quad m=\frac{1}{2}\ln{\bigg{(}\frac{\sigma_{T}^{2}}{x(x-1)}\bigg{)}},\quad\tau=\text{sgn}(\eta_{T})\mu_{T}-\frac{\sigma_{T}}{(x-1)^{\frac{1}{2}}},$
where sgn is the sign function and
[…]
With parameters $c,s,m,\tau$ from Lemma 4.2, the price of the basket option can be approximated by
$\widehat{C}(\widetilde{K}^{d},T):={{\mathbb{E}}}_{{{\mathbb{Q}}}}\big{[}\big{(}c(e^{sZ+m}+\tau)-\widetilde{K}^{d}\big{)}^{+}\big{]}$
or, more explicitly,
$\widehat{C}(\widetilde{K}^{d},T):={{\mathbb{E}}}_{{{\mathbb{Q}}}}\big{[}\big{(}e^{sZ+m}+\tau-\widetilde{K}^{d}\big{)}^{+}\big{]}$
if $\eta_{T}>0$ and
$\widehat{C}(\widetilde{K}^{d},T):={{\mathbb{E}}}_{{{\mathbb{Q}}}}\big{[}\big{(}-e^{sZ+m}-\tau-\widetilde{K}^{d}\big{)}^{+}\big{]}$
if $\eta_{T}<0$. The approximation for the price of the cross-currency basket call option is given by the following result, which is an immediate consequence of Theorem 2.3 in Hu et al. [10].
Lemma 4.3.
The following equalities hold
$\widehat{C}(\widetilde{K}^{d},T)=\begin{cases}e^{m+\frac{1}{2}s^{2}}+\tau-\widetilde{K}^{d},\quad&c=1,\,\widetilde{K}^{d}\leq\tau,\\ e^{m+\frac{1}{2}s^{2}}\Phi(d_{11})-(\widetilde{K}^{d}-\tau)\Phi(d_{12}),\quad&c=1,\,\widetilde{K}^{d}>\tau,\\ 0,\quad&c=-1,\,\widetilde{K}^{d}>-\tau,\\ -e^{m+\frac{1}{2}s^{2}}\Phi(d_{21})+(-\widetilde{K}^{d}-\tau)\Phi(d_{22}),\quad&c=-1,\,\widetilde{K}^{d}\leq-\tau,\end{cases}$
where the parameters $c,s,m,\tau$ are given by Lemma 4.2 and
$d_{11}$ $=\frac{-\ln(\widetilde{K}^{d}-\tau)+m+s^{2}}{s},\quad d_{12}=\frac{-\ln(\widetilde{K}^{d}-\tau)+m}{s},$
$d_{21}$ $=\frac{\ln(-\widetilde{K}^{d}-\tau)-m-s^{2}}{s},\quad d_{22}=\frac{\ln(-\widetilde{K}^{d}-\tau)-m}{s}.$
4.3 Superhedging of an Aggregated Effective Buffer EPS
It should be acknowledged there are still some weaknesses in the concept of the static hedging strategy for an aggregated effective EPS since it may be not easily available in practice. Option market makers are more likely to use a traded equity or market index as the underlying asset for options and options with the holder’s portfolio as the underlying asset can only be offered on OTC market. Hence, although Corollary 4.1 shows that, in theory, a perfect hedging of an EPS on a cross-currency reference portfolio is feasible, in reality, it may be difficult for the insurance company and EPS provider to implement the static hedging for an aggregated effective EPS. For that reason, we will consider the superhedging strategy for an aggregated effective EPS. Another option for the provider would be to built a stochastic model for the cross-currency market and use it to implement a dynamic hedging strategy based on actively traded assets and futures contracts on currencies.
Lemma 4.4.
The following inequalities hold, for every $(x,y)\in{\mathbb{R}}^{2}$,
$(x+y)^{+}\leq x^{+}+y^{+},$
$(x+y)^{+}\geq(x+y)\mathds{1}_{\{x\geq 0,\,y\geq 0\}}=x^{+}\mathds{1}_{\{y\geq 0\}}+y^{+}\mathds{1}_{\{x\geq 0\}},$
where the first inequality is strict on the set $A:=\{x<0,\,y>0\}\cup\{x>0,\,y<0\}$ and the second one is strict on the set $B:=\{x<0,\,x+y>0\}\cup\{y<0,\,x+y>0\}.$
Our goal is to introduce a superhedging strategy by using options with the underlying assets $S^{d}$ and $S^{f,e}$ introduced in Sections 3.3 and 3.4. Recall that the aggregate effective return satisfies $S^{w,e}_{T}=w\widehat{S}^{d}_{T}+(1-w)\widehat{S}^{f,e}_{T}$. Hence Lemma 4.4 leads to the following result.
Lemma 4.5.
The following inequalities hold, for every $w\in[0,1]$,
$\big{(}k^{l}_{i}-S^{w,e}_{T}\big{)}^{+}\leq w\big{(}k^{l}_{i}-\widehat{S}^{d}_{T}\big{)}^{+}+(1-w)\big{(}k^{l}_{i}-\widehat{S}^{f,e}_{T}\big{)}^{+},$
$\big{(}k^{l}_{i}-S^{w,e}_{T}\big{)}^{+}\geq w\big{(}k^{l}_{i}-\widehat{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\leq k^{l}_{i}\}}+(1-w)\big{(}k^{l}_{i}-\widehat{S}^{f,e}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq k^{l}_{i}\}},$
$\big{(}S^{w,e}_{T}-k^{g}_{j}\big{)}^{+}\leq w\big{(}\widehat{S}^{d}_{T}-k^{g}_{j}\big{)}^{+}+(1-w)\big{(}\widehat{S}^{f,e}_{T}-k^{g}_{j}\big{)}^{+},$
$\big{(}S^{w,e}_{T}-k^{g}_{j}\big{)}^{+}\geq w\big{(}\widehat{S}^{d}_{T}-k^{g}_{j}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\geq k^{g}_{j}\}}+(1-w)\big{(}\widehat{S}^{f,e}_{T}-k^{g}_{j}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq k^{g}_{j}\}}.$
Proof.   For the first two inequalities, we observe that
$k^{l}_{i}-S^{w,e}_{T}=w(k^{l}_{i}-\widehat{S}^{d}_{T})+(1-w)(k^{l}_{i}-\widehat{S}^{f,e}_{T})$
and we apply Lemma 4.4. For the last two, we use the equality
$S^{w,e}_{T}-k^{g}_{j}=w(\widehat{S}^{d}_{T}-k^{g}_{j})+(1-w)(\widehat{S}^{f,e}_{T}-k^{g}_{j})$
and, once again, we make use of Lemma 4.4. $\Box$
It is worth noting that the first inequality in Lemma 4.5 is strict on the event
$A^{l}_{i}:=\{\widehat{S}^{d}_{T}>k^{l}_{i},\,\widehat{S}^{f,e}_{T}<k^{l}_{i}\}\cup\{S^{d}_{T}<k^{l}_{i},\,\widehat{S}^{f,e}_{T}>k^{l}_{i}\}$
and the second one is strict on
$B^{l}_{i}:=\{\widehat{S}^{d}_{T}>k^{l}_{i},\,S^{w,e}_{T}>k^{l}_{i}\}\cup\{\widehat{S}^{f,e}_{T}>k^{l}_{i},\,S^{w,e}_{T}>k^{l}_{i}\}.$
The third one is strict on the event
$A^{g}_{j}:=\{\widehat{S}^{d}_{T}<k^{g}_{j},\,\widehat{S}^{f,e}_{T}>k^{g}_{j}\}\cup\{\widehat{S}^{d}_{T}>k^{g}_{j},\,\widehat{S}^{f,e}_{T}<k^{g}_{j}\}.$
and the last one is strict on
$B^{g}_{j}:=\{\widehat{S}^{d}_{T}<k^{g}_{j},\,S^{w,e}_{T}>k^{g}_{j}\}\cup\{\widehat{S}^{f,e}_{T}<k^{g}_{j},\,S^{w,e}_{T}>k^{g}_{j}\}.$
Lemma 4.5 can be used to construct a particular superhedging strategy for an aggregated effective buffer EPS. From Corollary 4.1, we obtain for the buffer EPS
$\mbox{\rm EPS\,}^{w,e}_{T}=N\big{(}p_{2}\widehat{P}_{T}(S^{w,e},k^{l}_{1},T)-f_{2}\widehat{C}_{T}(S^{w,e},k^{g}_{1},T)\big{)}$
where $k^{l}_{1}=1+l_{1}$ and $k^{g}_{1}=1+g_{1}$. Hence it suffices to set $N=1$ and consider the contingent claim
$Y:=\psi_{B}(S^{w,e}_{T})=p_{2}\big{(}k^{l}_{1}-S^{w,e}_{T}\big{)}^{+}-f_{2}\big{(}S^{w,e}_{T}-k^{g}_{1}\big{)}^{+}$
where $p_{2}$ and $f_{2}$ are strictly positive numbers. We wish to find a contingent claim $\widehat{Y}$ such that $\widehat{Y}\geq Y$ and static hedging of $\widehat{Y}$ is feasible with traded options. By combining the first and the last inequality in Lemma 4.5, we obtain
$Y$ $=p_{2}\big{(}k^{l}_{1}-S^{w,e}_{T}\big{)}^{+}-f_{2}\big{(}S^{w,e}_{T}-k^{g}_{1}\big{)}^{+}\leq p_{2}w\big{(}k^{l}_{1}-\widehat{S}^{d}_{T}\big{)}^{+}-f_{2}w\big{(}\widehat{S}^{d}_{T}-k^{g}_{1}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}>k^{g}_{1}\}}$
$+p_{2}(1-w)\big{(}k^{l}_{1}-\widehat{S}^{f,e}_{T}\big{)}^{+}-f_{2}(1-w)\big{(}\widehat{S}^{f,e}_{T}-k^{g}_{1}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}>k^{g}_{1}\}}=:\widehat{Y}.$
Since $\widehat{Y}$ dominates $Y$ it is clear that a replicating strategy for $\widehat{Y}$ is also a superhedging strategy for $Y$ and, in fact, it is a strict superhedging of $Y$ on the event $A^{l}_{1}\cup B^{g}_{1}$. Let us denote such a strategy by $H^{w,e,s,b}$ so that $H_{T}^{w,e,s,b}=\widehat{Y}$. Since our goal is to obtain a static trading strategy, we introduce European call options with the payoff in the domestic currency given by
$C^{*}_{T}(\widehat{S}^{d},K,T,\widehat{S}^{f,e}):=\big{(}\widehat{S}^{d}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\geq K\}}$
and
$C^{*}_{T}(\widehat{S}^{f,e},K,T,\widehat{S}^{d}):=\big{(}\widehat{S}^{f,e}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq K\}}.$
Then we have the following corollary to Lemma 4.5.
Corollary 4.2.
Let $H^{w,e,s,b}$ be a static superhedging hedging strategy for an aggregated effective buffer EPS such that $H_{T}^{w,e,s,b}=\widehat{Y}$. Then the initial cost of $H^{w,e,s,b}$ equals
$H^{w,e,s,b}_{0}$ $=w\Big{(}p_{2}P_{0}(\widehat{S}^{d},k^{l}_{1},T)-f_{2}C^{*}_{0}(\widehat{S}^{d},k^{g}_{1},T,\widehat{S}^{f,e})\Big{)}$
$+(1-w)\Big{(}p_{2}\widetilde{P}_{0}(\widehat{S}^{f,e},k^{l}_{1},T)-f_{2}C^{*}_{0}(\widehat{S}^{f,e},k^{g}_{1},T,\widehat{S}^{d})\Big{)}.$
To find the superhedging cost $H^{w,e,s,b}_{0}$, it suffices to compute the prices $C^{*}_{0}(\widehat{S}^{d},k^{g}_{1},T,\widehat{S}^{f,e})$ and $C^{*}_{0}(\widehat{S}^{f,e},k^{g}_{1},T,\widehat{S}^{d})$. Let $\Phi_{2}(x_{1},x_{2},\rho)$ denote the c.d.f. of the two-dimensional standard Gaussian distribution with the correlation coefficient $\rho$. We set
$\rho:=\frac{(\bm{\sigma}^{f}+\bm{\sigma}^{q})(\bm{\sigma}^{d})^{*}}{\|\bm{\sigma}^{d}\|\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|}.$
Proposition 4.2.
The following equality holds, for every $K>0$,
$C^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f,e})=\Phi_{2}(g_{1},g_{2},\rho)-\widetilde{K}\Phi_{2}(d_{1},d_{2},\rho)$
where $\widetilde{K}=e^{-r^{d}T}K$ and
$d_{1}$ $=\frac{-\ln\widetilde{K}-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}{\|\bm{\sigma}^{d}\|\sqrt{T}},\quad d_{2}=\frac{-\ln\widetilde{K}-\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2}T}{\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|\sqrt{T}},$
$g_{1}$ $=\frac{-\ln\widetilde{K}+\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}{\|\bm{\sigma}^{d}\|\sqrt{T}},\quad g_{2}=\frac{-\ln\widetilde{K}+(\bm{\sigma}^{f}+\bm{\sigma}^{q})(\bm{\sigma}^{d})^{*}T-\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2}T}{\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|\sqrt{T}}.$
Proof.   Recall that $\widehat{S}^{d}$ and $\widehat{S}^{f,e}$ satisfy under the domestic martingale measure ${{\mathbb{Q}}}^{d}$
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}r^{d}\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}\big{)},$
$d\widehat{S}^{f,e}_{t}=\widehat{S}^{f,e}_{t}\big{(}r^{d}\,dt+(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\textbf{W}_{t}^{d}\big{)},$
with initial conditions $\widehat{S}^{d}_{0}=\widehat{S}^{f,e}_{0}=1$. We observe that
$C^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f,e})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widehat{S}^{d}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\geq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{d}_{T}-\widetilde{K}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\geq K\}}\big{]}$
$={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{d}_{T}-\widetilde{K}\big{)}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{d}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\}}\big{]}-\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\big{]}=:I_{1}-\widetilde{K}I_{2}$
where $\widetilde{S}^{d}_{t}:=e^{-r^{d}t}\widehat{S}^{d}_{t}$ so that $d\widetilde{S}^{d}_{t}=\widetilde{S}^{d}_{t}\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}$ and $\widetilde{K}:=e^{-r^{d}T}K$. We introduce a probability measure ~$\widetilde{{\mathbb{Q}}}^{d}$ equivalent to ${{\mathbb{Q}}}^{d}$ on $(\Omega,\mathcal{F}_{T})$ by the Radon-Nikodym density
$\frac{d\widetilde{{\mathbb{Q}}}^{d}}{d{{\mathbb{Q}}}^{d}}:=e^{\bm{\sigma}^{d}\,\textbf{W}_{T}^{d}-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}.$ (8)
Then the process $\widetilde{\textbf{W}}_{t}^{d},\,t\in[0,T]$, which is given by the equality $\widetilde{\textbf{W}}_{t}^{d}=\textbf{W}_{t}^{d}-(\bm{\sigma}^{d})^{*}t$ is a standard Brownian motion under ~$\widetilde{{\mathbb{Q}}}^{d}$ and the dynamics of $\widehat{S}^{d}$ and $\widehat{S}^{f,e}$ under ~$\widetilde{{\mathbb{Q}}}^{d}$ are
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}(r^{d}+\|\bm{\sigma}^{d}\|^{2})\,dt+\bm{\sigma}^{d}\,d\widetilde{\textbf{W}}_{t}^{d}\big{)},$
$d\widehat{S}^{f,e}_{t}=\widehat{S}^{f,e}_{t}\big{(}\big{(}r^{d}+(\bm{\sigma}^{f}+\bm{\sigma}^{q})(\bm{\sigma}^{d})^{*}\big{)}\,dt+(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\widetilde{\textbf{W}}_{t}^{d}\big{)}.$
It is now easy to check that
$I_{1}=\widetilde{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\big{]}=\Phi_{2}(g_{1},g_{2},\rho),$
$I_{2}={{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\big{]}=\Phi_{2}(d_{1},d_{2},\rho),$
where $g_{1},g_{2},d_{1}$ and $d_{2}$ are given in the statement of the lemma. $\Box$
The result for $C^{*}_{0}(\widehat{S}^{f,e},K,T,\widehat{S}^{d})$ is analogous though, of course, it requires minor adjustments.
Proposition 4.3.
The following equality holds, for every $K>0$,
$C^{*}_{0}(\widehat{S}^{f,e},K,T,\widehat{S}^{d})=\Phi_{2}(h_{1},h_{2},\rho)-\widetilde{K}\Phi_{2}(d_{1},d_{2},\rho)$
where $d_{1}$ and $d_{2}$ are given in Propositon 4.2 and
$h_{1}=\frac{-\ln\widetilde{K}+(\bm{\sigma}^{f}+\bm{\sigma}^{q})(\bm{\sigma}^{d})^{*}T-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}{\|\bm{\sigma}^{d}\|\sqrt{T}},\quad h_{2}=\frac{-\ln\widetilde{K}+\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2}T}{\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|\sqrt{T}}.$
Proof.   We have
$C^{*}_{0}(\widehat{S}^{f,e},K,T,\widehat{S}^{d})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widehat{S}^{f,e}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{f,e}_{T}-\widetilde{K}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq K\}}\big{]}$
$={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{f,e}_{T}-\widetilde{K}\big{)}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{f,e}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\}}\big{]}-\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f,e}_{T}>K\big{]}=:\widehat{I}_{1}-\widetilde{K}I_{2}.$
We introduce an equivalent probability measure ^$\widehat{{\mathbb{Q}}}^{d}$ on $(\Omega,\mathcal{F}_{T})$ by the Radon-Nikodym density
$\frac{d\widehat{{\mathbb{Q}}}^{d}}{d{{\mathbb{Q}}}^{d}}:=e^{(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,\textbf{W}_{T}^{d}-\frac{1}{2}\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2}T}.$ (9)
Then the process $\widehat{\textbf{W}}_{t}^{d},\,t\in[0,T]$ given by the equality $\widehat{\textbf{W}}_{t}^{d}=\textbf{W}_{t}^{d}-(\bm{\sigma}^{f}+\bm{\sigma}^{q})^{*}t$ is a standard Brownian motion under ^$\widehat{{\mathbb{Q}}}^{d}$ and the dynamics of $\widehat{S}^{d}$ and $\widehat{S}^{f,e}$ under ^$\widehat{{\mathbb{Q}}}^{d}$ are
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}r^{d}\,dt+(\bm{\sigma}^{f}+\bm{\sigma}^{q})(\bm{\sigma}^{d})^{*}\,dt+\bm{\sigma}^{d}\,d\widehat{\textbf{W}}_{t}^{d}\big{)},$
$d\widehat{S}^{f,e}_{t}=\widehat{S}^{f,e}_{t}\big{(}(r^{d}+\|\bm{\sigma}^{f}+\bm{\sigma}^{q}\|^{2})\,dt+(\bm{\sigma}^{f}+\bm{\sigma}^{q})\,d\widehat{\textbf{W}}_{t}^{d}\big{)}.$
Hence it is easy to check that the stated expressions are valid. $\Box$
4.4 Superhedging of an Aggregated Effective Floor EPS
Our next goal is to study a static superhedging strategy for an aggregated effective floor EPS. For $N=1$, we now have that
$\mbox{\rm EPS\,}^{w,e}_{T}=\psi_{F}(S^{w,e}_{T})=-p_{1}\,\widehat{P}_{T}(S^{w,e},k^{l}_{1},T)+p_{1}\widehat{P}_{T}(S^{w,e},1,T)-f_{2}\,\widehat{C}_{t}(S^{w,e},k^{g}_{1},T)$
and we denote
$P^{*}_{T}(\widehat{S}^{d},K,T,\widehat{S}^{f,e}):=\big{(}K-\widehat{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\leq K\}}$
and
$P^{*}_{T}(\widehat{S}^{f,e},K,T,\widehat{S}^{d}):=\big{(}K-\widehat{S}^{f,e}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq K\}}.$
The following corollary to Lemma 4.5 gives the initial cost of a static superhedging strategy for an aggregated effective floor EPS.
Corollary 4.3.
The initial cost of a static superhedging strategy $H^{w,e,s,f}$ for an aggregated effective floor EPS with $N=1$ satisfies
$H^{w,e,s,f}_{0}$ $=w\Big{(}p_{1}P_{0}(\widehat{S}^{d},1,T)-p_{1}P^{*}_{0}(\widehat{S}^{d},k^{l}_{1},T)-f_{2}C^{*}_{0}(\widehat{S}^{d},k^{g}_{1},T,\widehat{S}^{f,e})\Big{)}$
$+(1-w)\Big{(}p_{1}\widetilde{P}_{0}(\widehat{S}^{f,e},1,T)-f_{2}C^{*}_{0}(\widehat{S}^{f,e},k^{g}_{1},T,\widehat{S}^{d})-p_{1}P^{*}_{0}(\widehat{S}^{f,e},k^{l}_{1},T,\widehat{S}^{d})\Big{)}.$
To find an explicit representation for the cost $H^{w,e,s,f}_{0}$ it suffices to combine Propositions 4.2 and 4.3 with the following result.
Proposition 4.4.
The following equalities are valid, for every $K>0$,
$P^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f,e})=\widetilde{K}\Phi_{2}(-d_{1},-d_{2},\rho)-\Phi_{2}(-g_{1},-g_{2},\rho),$
$P^{*}_{0}(\widehat{S}^{f,e},K,T,\widehat{S}^{d})=\widetilde{K}\Phi_{2}(-d_{1},-d_{2},\rho)-\Phi_{2}(-h_{1},-h_{2},\rho).$
Proof.   We first compute
$P^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f,e})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}K-\widehat{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\leq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{K}-\widetilde{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f,e}_{T}\leq K\}}\big{]}$
$=\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f,e}_{T}<K\big{]}-{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{d}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f,e}_{T}<K\}}\big{]}.$
To obtain the first equality we use the change of a probability measure from the proof of Proposition 4.2. For the second equality, we observe that
$P^{*}_{0}(\widehat{S}^{f,e},K,T,\widehat{S}^{d})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}K-\widehat{S}^{f,e}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq K\}}\big{]}$
$=\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f,e}_{T}<K\big{]}-{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{f,e}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f,e}_{T}<K\}}\big{]}.$
and we use the change of a probability measure introduced in the proof of Proposition 4.3. $\Box$
4.5 Approximate Hedging of an Aggregated Effective Buffer EPS
The superhedging strategy $H^{w,e,s}$ combined with the premium $\widehat{c}$ ensures that the provider will never lose money and thus it can be seen as a good choice for the provider. However, the value $\widehat{c}$ is not necessarily competitive and may be too high for the buyer of the EPS since it does not represent an arbitrage-free price. It thus make sense to search for a less expensive hedging strategy even if it may generate a loss in some scenarios. For lack of a better term, we will refer to that situation as an approximate hedging. Of course, we would like to find an approximate hedging with some desirable properties, such as closeness to superhedging and convenience of implementation in practice.
If the indicator functions in definition of $\widehat{Y}$ are omitted, then we obtain the claim
$\widetilde{Y}=w[p_{2}\big{(}k^{l}_{1}-S^{d}_{T}\big{)}^{+}-f_{2}\big{(}k^{g}_{1}-S^{d}_{T}\big{)}^{+}]+(1-w)[p_{2}\big{(}k^{l}_{1}-S^{f,e}_{T}\big{)}^{+}-f_{2}\big{(}k^{g}_{1}-S^{f,e}_{T}\big{)}^{+}]$
which coincides with $\widehat{Y}$ on the complement of the event $B^{g}_{1}$. On the one hand, it should be acknowledged that the claim $\widetilde{Y}$ does not dominate $Y$, in general. On the other hand, to replicate $\widetilde{Y}$ it suffices to use European options with the underlying assets $S^{d}$ and $S^{f,e}$ to build the hedging portfolios from Propositions 3.1 and 3.3. Then an approximate replication is obtained by combining the domestic and effective foreign static hedging strategies with weights $w$ and $1-w$, respectively. Formally, the wealth $H^{w,e,a}$ of the approximate hedging strategy for the aggregated effective buffer EPS satisfies
$H^{w,e,a}_{T}=w\widehat{H}^{d}_{T}+(1-w)\widehat{H}^{f,e}_{T}$
where $\widehat{H}^{d}_{T}$ and $\widehat{H}^{f,e}_{T}$ are given by Corollaries 3.1 and 3.3, respectively. The strategy $H^{w,e,a}$ does not satisfy the superhedging condition $H^{w,e,a}_{T}\geq\widehat{H}^{w,e}_{T}$ and thus it may result in provider’s losses in some adverse market scenarios. However, it only requires to use on simple options on reference portfolios $S^{d}$ and $S^{f}$.
Corollary 4.4.
The initial cost of the approximate hedging strategy $H^{w,e,a}$ for an effective EPS with $N=1$ equals
$H^{w,e,a}_{0}=w\widehat{H}^{d}_{0}+(1-w)\widehat{H}^{f,e}_{0}.$
Notice that the static hedging strategy from Corollary 4.2 has a similar structure to two separate hedging strategies from Propositions 3.1 and 3.3. The main differences between the strategy from Corollary 4.2 and separate strategies are the underlying assets and the strikes of European options used in hedging portfolios. In separate hedging strategies, the provider considers domestic equity and foreign equity independently and thus the strikes of options are only related to each economy and, as a consequence, thus this also applies to the combined superhedging strategy. In contrast, a static hedging strategy given by $\widehat{H}^{w,e}_{T}$ hinges on the cross-currency reference portfolio as an underlying asset and thus strikes are based on the whole portfolio.
5 Pricing and Hedging of an Aggregated Quanto EPS
Let us assume that the exchange rate $\overline{Q}$ has been fixed at time 0 and at any date $t>0$ the domestic value of foreign holdings is calculated using the exchange rate $\overline{Q}$, which in fact is an investment opportunity offered to investors by some financial companies. In principle, the exchange rate $\overline{Q}$ in a quanto contract is completely arbitrary but we always set $\overline{Q}=Q_{0}$ to simplify the presentation. Then the portfolio’s quanto wealth calculated using the exchange rate $Q_{0}$ equals, for every $t\in[0,T]$,
$X^{w,q}_{t}:=\alpha_{0}S^{d}_{t}+\beta_{0}Q_{0}S^{f}_{t}=N\big{[}wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f}_{t}(S^{f}_{0})^{-1}\big{]}=NS^{w,q}_{t}$
where $S^{w,q}_{t}:=w(1+R^{d}_{t})+(1-w)(1+R^{f}_{t})$ with $S^{w,q}_{0}=1$ is a counterpart of the process $S^{w,e}$ introduced in Section 4. The corresponding rate of return on $S^{w,q}$ is called aggregated quanto return and denoted by $R^{w,q}$. For every $N>0$ and $w\in[0,1]$, the aggregated quanto return $R^{w,q}$ on a cross-currency portfolio equals, for every $t\in[0,T]$,
$R^{w,q}_{t}:=\frac{X^{w,q}_{t}-X^{w,q}_{0}}{X^{w,q}_{0}}=S^{w,q}_{t}-1=wR^{d}_{t}+(1-w)R^{f}_{t}.$
Definition 5.1.
The provider’s cash flow at time $T$ in the aggregated quanto EPS is given in the domestic currency and equals
$\mbox{\rm EPS\,}^{w,q}_{T}:=N\psi_{a}(R^{w,q}_{T})=N\psi_{a}\big{(}wR^{d}_{T}+(1-w)R^{f}_{T}\big{)}.$ (10)
It is natural to focus on the process $S^{w,q}$ whose value in the domestic currency equals, for every $t\in[0,T]$,
$S^{w,q}_{t}=wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f}_{t}(S^{f}_{0})^{-1}.$
Similar to Section 3.5, we write $\overline{C}_{T}(S^{w,q},K^{d},T):=(S^{w,q}_{T}-K^{d})^{+}$ and $\overline{P}_{T}(S^{w,q},K^{d},T):=(K^{d}-S^{w,q}_{T})^{+}$ where the strike $K^{d}$ is given in the domestic currency and, similarly, we denote by $\overline{C}(S^{w,q},K^{d},T)$ and $\overline{P}(S^{w,q},K^{d},T)$ their respective price processes. Then a static hedging strategy for the cross-currency quanto EPS is given by the following result.
Corollary 5.1.
The static hedging in terms of options written on the process $S^{w,q}$ for the aggregated quanto EPS is given by Proposition 2.1 with $\psi=\psi_{a}$ and $X=S^{w,q}$, that is, $\mbox{\rm EPS\,}^{w,q}_{T}=\widehat{H}^{w,q}_{T}$ where, for every $t\in[0,T]$,
$\widehat{H}^{w,q}_{T}=N\sum_{i=0}^{n}(p_{i+1}-p_{i})\,\overline{P}_{t}(S^{w,q},k^{l}_{i},T)-N\sum_{j=0}^{m}(f_{j+1}-f_{j})\,\overline{C}_{t}(S^{w,q},k^{g}_{j},T)$
where $k^{l}_{i}=1+l_{i}$ and $k^{g}_{j}=1+g_{j}$ for every $i=0,1,\dots,n$ and $j=0,1,\dots,m$.
5.1 Approximate Pricing using Geometric Averaging Method
To find approximate values for prices of the quanto basket call and put options, and hence also the approximate price of an aggregated quanto EPS, we follow the approach from Section 4.1 with suitable minor adjustments. We now have
$S^{w,q}_{t}=wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f}_{t}(S^{f}_{0})^{-1}=w\widehat{S}^{d}_{t}+(1-w)\widehat{S}^{f}_{t}$
where, for the sake of computational convenience, we set $\widehat{S}^{d}_{t}:=S^{d}_{t}(S^{d}_{0})^{-1}$ and $\widehat{S}^{f}_{t}:=S^{f}_{t}(S^{f}_{0})^{-1}$ so that $\widehat{S}^{d}_{0}=\widehat{S}^{f}_{0}=1$ and
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}r^{d}\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}\big{)},$
$d\widehat{S}^{f}_{t}=\widehat{S}^{f}_{t}\big{(}(r^{f}-\bm{\sigma}^{f}(\bm{\sigma}^{q})^{*})\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}^{d}\big{)}.$
We wish to compute the price of the quanto basket call option
$\overline{C}_{0}(S^{w,q},K^{d},T)=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}(S^{w,q}_{T}-K^{d})^{+}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}w\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f}_{T}-\widetilde{K}^{d}\big{)}^{+}\big{]}$
where we denote $\widetilde{S}^{d}_{t}:=e^{-r^{d}t}\widehat{S}^{d}_{t}$ and $\widetilde{S}^{f}_{t}:=e^{-r^{d}t}\widehat{S}^{f}_{t}$ and $\widetilde{K}^{d}:=e^{-r^{d}T}K^{d}$ so that
$d\widetilde{S}^{d}_{t}=\widetilde{S}^{d}_{t}\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d},$
$d\widetilde{S}^{f}_{t}=\widetilde{S}^{f}_{t}\big{(}\delta^{q}\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}^{d}\big{)},$
where $\delta^{q}:=r^{f}-r^{d}-\bm{\sigma}^{f}(\bm{\sigma}^{q})^{*}$ and $\widetilde{S}^{d}_{0}=\widetilde{S}^{f,e}_{0}=1$. We set
$\bar{G}_{T}:=\big{(}\widetilde{S}^{d}_{T}\big{)}^{w}\big{(}\widetilde{S}^{f}_{T}\big{)}^{(1-w)}=e^{(w\bm{\sigma}^{d}+(1-w)\bm{\sigma}^{f})\textbf{W}_{T}^{d}-\frac{1}{2}\bar{\gamma}T}=e^{\bar{\xi}_{T}-\frac{1}{2}\bar{\gamma}T}$
where $\bar{\gamma}:=w\|\sigma^{d}\|^{2}+(1-w)\|\sigma^{f}\|^{2}-2(1-w)\delta^{q}$ and $\bar{\xi}_{T}$ is a Gaussian random variable under ${{\mathbb{Q}}}^{d}$ with null expected value and the variance $\bar{v}^{2}T$ where
$\bar{v}^{2}:=\|w\bm{\sigma}^{d}+(1-w)\bm{\sigma}^{f}\|^{2}=w^{2}\|\sigma^{d}\|^{2}+2w(1-w)\sigma^{d}(\sigma^{f})^{*}+(1-w)^{2}\|\sigma^{f}\|^{2}.$
Then the random variable $\bar{G}_{T}$ has the lognormal distribution under ${{\mathbb{Q}}}^{d}$ with the expected value
${{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}(\bar{G}_{T})$ $=e^{\frac{1}{2}(\bar{v}^{2}-\bar{\gamma})T}=\exp\Big{(}-\frac{1}{2}w(1-w)\big{(}\|\sigma^{d}\|^{2}-2\sigma^{d}(\sigma^{f})^{*}+\|\sigma^{f}\|^{2}\big{)}T+(1-w)\delta^{q}T\Big{)}$
$=\exp\Big{(}-\frac{1}{2}w(1-w)\|\sigma^{d}-\sigma^{f}\|^{2}T+(1-w)\delta^{q}T\Big{)}=:\bar{\lambda}.$
Similar to Section 4.1, we contend that a plausible approximation for the price of the quanto basket call option is given by the expected value
$C^{\bar{G}}_{0}(\bar{K}^{d},T)={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{(}\big{(}\bar{G}_{T}-\bar{K}^{d}\big{)}^{+}\big{)}$
where $\bar{K}^{d}$ equals
$\bar{K}^{d}:=\widetilde{K}^{d}+{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\Big{(}\bar{G}_{T}-\big{(}w_{t}\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f}_{T}\big{)}\Big{)}=\widetilde{K}^{d}+\bar{\lambda}-\kappa$
where
$\kappa:={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{(}w\widetilde{S}^{d}_{T}+(1-w)\widetilde{S}^{f}_{T}\big{)}=w\widetilde{S}^{d}_{0}+(1-w)e^{\delta^{q}T}\widetilde{S}^{f}_{0}=w+(1-w)e^{\delta^{q}T}$
where the last equality follows from the dynamics of $\widetilde{S}^{d}$ and $\widetilde{S}^{f,e}$ under ${{\mathbb{Q}}}^{d}$. Elementary computations now yield the following result, which is a counterpart of Proposition 4.1.
Proposition 5.1.
The price $\overline{C}_{0}(S^{w,q},K^{d},T)$ of the quanto basket call option with the payoff $\overline{C}_{T}(S^{w,e},K^{d},T)=(S^{w,e}_{T}-K^{d})^{+}$ can be approximated by
$C^{\bar{G}}_{0}(\bar{K}^{d},T)=\bar{\lambda}\Phi\big{(}\bar{d}_{+}(\bar{K}^{d},T)\big{)}-\bar{K}^{d}\Phi\big{(}\bar{d}_{-}(\bar{K}^{d},T)\big{)}$
where $\bar{d}_{+}(\bar{K}^{d},T)$ and $\bar{d}_{-}(\bar{K}^{d},T)$ are given by
$\bar{d}_{\pm}(\bar{K}^{d},T)=\frac{1}{\bar{v}\sqrt{T}}\Big{[}\ln\frac{\bar{\lambda}}{\bar{K}^{d}}\pm\frac{1}{2}\bar{v}^{2}T\Big{]}.$
The price $\overline{P}_{0}(S^{w,q},K,T)$ of the quanto basket put option with the payoff $\overline{P}_{T}(S^{w,e},K^{d},T)=(K^{d}-S^{w,e}_{T})^{+}$ can be approximated by
$P^{\bar{G}}_{0}(\bar{K}^{d},T)=\bar{K}^{d}\Phi\big{(}-\bar{d}_{-}(\bar{K}^{d},T)\big{)}-\bar{\lambda}\Phi\big{(}-\bar{d}_{+}(\widehat{K}^{d},T)\big{)}.$
Remark 5.1.
The put-call parity for quanto basket options reads
$\overline{C}_{0}(S^{w,e},K^{d},T)-\overline{P}_{0}(S^{w,e},K^{d},T)=w+(1-w)e^{\delta^{q}T}-\widetilde{K}^{d}$
whereas for the approximated values it can be computed as follows
$C^{G}_{0}(\widehat{K}^{d},T)-P^{G}_{0}(\widehat{K}^{d},T)=\bar{\lambda}-\widehat{K}^{d}=\bar{\lambda}-(\widetilde{K}^{d}+\bar{\lambda}-\kappa)=\kappa-\widetilde{K}^{d}=w+(1-w)e^{\delta^{q}T}-\widetilde{K}^{d}.$
5.2 Approximate Pricing using Moment Matching Method
It is clear that moment matching method can also be used for approximate pricing of quanto basket options. To this end, it suffices to compute the first three moments of $\widetilde{S}^{w,q}_{T}$ and then identify the parameters $\bar{c},\bar{s},\bar{m}$ and $\bar{\tau}$ by solving the moment matching conditions $\overline{M}_{i}=M_{i}(\bar{c},\bar{s},\bar{m},\bar{\tau})$ for $i=1,2,3$ where $M_{i}$ is given by (4.2) and $\overline{M}_{i}$ is given by Lemma 4.5. Then parameters $\bar{c},\bar{s},\bar{m}$ and $\bar{\tau}$ can be computed using explicit formulae from Lemma 4.2 and an approximation for the price of an aggregated quanto EPS can be obtained using Lemma 4.3 with the modified set of parameters.
Recall that the dynamics of the processes $\widetilde{S}^{d}$ and $\widetilde{S}^{f}$ under ${\mathbb{Q}}^{d}$ are
$d\widetilde{S}^{d}_{t}=\widetilde{S}^{d}_{t}\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d},$
$d\widetilde{S}^{f}_{t}=\widetilde{S}^{f}_{t}\big{(}\delta^{q}\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}^{d}\big{)},$
with initial conditions $\widetilde{S}^{d}_{0}=\widetilde{S}^{f,e}_{0}=1$. Consequently, for the process $\widetilde{S}^{w,q}$ given by the equality $\widetilde{S}^{w,q}_{t}=w\widetilde{S}^{d}_{t}+(1-w)\widetilde{S}^{f}_{t}$ we have that
$\widetilde{S}^{w,q}_{T}=we^{\bm{\sigma}^{d}\textbf{W}_{T}^{d}-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}+(1-w)e^{\bm{\sigma}^{f}\textbf{W}_{T}^{d}+\delta^{q}T-\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}T}.$
The initial price of the aggregated quanto call equals
$\overline{C}_{0}(S^{w,q},K^{d},T)={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{(}\big{(}\widetilde{S}^{w,q}_{T}-\widetilde{K}^{d}\big{)}^{+}\big{)}$
and thus it suffices to use a straightforward modification of Lemma 4.1.
Lemma 5.1.
Let $w_{1}=w,\,w_{2}=(1-w)e^{\delta^{q}T},\,\bm{\sigma}_{1}=\bm{\sigma}^{d}$ and $\bm{\sigma}_{2}=\bm{\sigma}^{f}$. Then the first three moments of the random variable $\widetilde{S}^{w,q}_{T}$ under the martingale measure ${\mathbb{Q}}^{d}$ are given by the following expressions
$\overline{M}_{1}$ $=w_{1}+w_{2}=w+(1-w)e^{\delta^{q}T},$
$\overline{M}_{2}$ $=w_{1}^{2}e^{\|\bm{\sigma}_{1}\|^{2}T}+2w_{1}w_{2}e^{\bm{\sigma}_{1}\bm{\sigma}_{2}^{*}T}+w_{2}^{2}e^{\|\bm{\sigma}_{2}\|^{2}T},$
$\overline{M}_{2}$ $=w_{1}^{3}e^{3\|\bm{\sigma}_{1}\|^{2}T}+3w_{1}^{2}w_{2}e^{\big{(}\|\bm{\sigma}_{1}\|^{2}+2\bm{\sigma}_{1}\bm{\sigma}_{2}^{*}\big{)}T}+3w_{1}w_{2}^{2}e^{\big{(}\|\bm{\sigma}_{2}\|^{2}+2\bm{\sigma}_{1}\bm{\sigma}_{2}^{*}\big{)}T}+w_{2}^{3}e^{3\|\bm{\sigma}_{2}\|^{2}T}.$
Then Lemma 4.3 with parameters $\bar{c},\bar{s},\bar{m}$ and $\bar{\tau}$ satisfying the moment matching conditions $\overline{M}_{i}=M_{i}(\bar{c},\bar{s},\bar{m},\bar{\tau})$ yields the approximate prices of quanto basket call and put options.
5.3 Superhedging of an Aggregated Quanto Buffer EPS
We will now analyse a static superhedging strategy for an aggregated quanto buffer EPS. Recall that the process $S^{w,q}$ equals, for every $t\in[0,T]$,
$S^{w,q}_{t}=wS^{d}_{t}(S^{d}_{0})^{-1}+(1-w)S^{f}_{t}(S^{f}_{0})^{-1}=w\widehat{S}^{d}_{t}+(1-w)\widehat{S}^{f}_{t}.$
Similar to Lemma 4.5, we now have the following inequalities, for every $w\in[0,1]$,
$\big{(}k^{l}_{i}-S^{w,q}_{T}\big{)}^{+}\leq w\big{(}k^{l}_{i}-\widehat{S}^{d}_{T}\big{)}^{+}+(1-w)\big{(}k^{l}_{i}-\widehat{S}^{f}_{T}\big{)}^{+},$
$\big{(}k^{l}_{i}-S^{w,q}_{T}\big{)}^{+}\geq w\big{(}k^{l}_{i}-\widehat{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\leq k^{l}_{i}\}}+(1-w)\big{(}k^{l}_{i}-\widehat{S}^{f}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq k^{l}_{i}\}},$
$\big{(}S^{w,q}_{T}-k^{g}_{j}\big{)}^{+}\leq w\big{(}\widehat{S}^{d}_{T}-k^{g}_{j}\big{)}^{+}+(1-w)\big{(}\widehat{S}^{f}_{T}-k^{g}_{j}\big{)}^{+},$
$\big{(}S^{w,q}_{T}-k^{g}_{j}\big{)}^{+}\geq w\big{(}\widehat{S}^{d}_{T}-k^{g}_{j}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\geq k^{g}_{j}\}}+(1-w)\big{(}\widehat{S}^{f}_{T}-k^{g}_{j}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq k^{g}_{j}\}}.$
From Corollary 4.1 with $S^{w,e}$ substituted with $S^{w,q}$, we obtain for the aggregated quanto buffer EPS with the notional principal $N=1$
$\mbox{\rm EPS\,}^{w,q}_{T}=\psi_{B}(S^{w,q}_{T})=p_{2}\overline{P}_{T}(S^{w,q},k^{l}_{1},T)-f_{2}\overline{C}_{T}(S^{w,q},k^{g}_{1},T)$
where $k^{l}_{1}=1+l_{1}$ and $k^{g}_{1}=1+g_{1}$ and thus
$\mbox{\rm EPS\,}^{w,q}_{T}$ $=p_{2}\big{(}k^{l}_{1}-S^{w,q}_{T}\big{)}^{+}-f_{2}\big{(}S^{w,q}_{T}-k^{g}_{1}\big{)}^{+}\leq p_{2}w\big{(}k^{l}_{1}-\widehat{S}^{d}_{T}\big{)}^{+}-f_{2}w\big{(}\widehat{S}^{d}_{T}-k^{g}_{1}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}>k^{g}_{1}\}}$
$+p_{2}(1-w)\big{(}k^{l}_{1}-\widehat{S}^{f}_{T}\big{)}^{+}-f_{2}(1-w)\big{(}\widehat{S}^{f}_{T}-k^{g}_{1}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}>k^{g}_{1}\}}=:\overline{Y}$
where the random variable $\overline{Y}$ manifestly dominates $\mbox{\rm EPS\,}^{w,q}_{T}$. We now denote
$\overline{C}^{*}_{T}(\widehat{S}^{d},K,T,\widehat{S}^{f}):=\big{(}\widehat{S}^{d}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\geq K\}}$
and
$\overline{C}^{*}_{T}(\widehat{S}^{f},K,T,\widehat{S}^{d}):=\big{(}\widehat{S}^{f}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq K\}}$
and we proceed as in Section 4.3.
Corollary 5.2.
Let $H^{w,q,s,b}$ be a static superhedging strategy for an aggregated quanto buffer EPS such that $H_{T}^{w,q,s,b}=\overline{Y}$. Then the initial cost of $H^{w,q,s,b}$ equals
$H^{w,q,s,b}_{0}$ $=w\Big{(}p_{2}P_{0}(\widehat{S}^{d},k^{l}_{1},T)-f_{2}\overline{C}^{*}_{0}(\widehat{S}^{d},k^{g}_{1},T,\widehat{S}^{f})\Big{)}$
$+(1-w)\Big{(}p_{2}\overline{P}_{0}(\widehat{S}^{f},k^{l}_{1},T)-f_{2}\overline{C}^{*}_{0}(\widehat{S}^{f},k^{g}_{1},T,\widehat{S}^{d})\Big{)}.$
To find the superhedging cost $H^{w,q,s,b}_{0}$, it now suffices to compute the prices $\overline{C}^{*}_{0}(\widehat{S}^{d},k^{g}_{1},T,\widehat{S}^{f,e})$ and $\overline{C}^{*}_{0}(\widehat{S}^{f},k^{g}_{1},T,\widehat{S}^{d})$. We now set
$\bar{\rho}:=\frac{\bm{\sigma}^{f}(\bm{\sigma}^{d})^{*}}{\|\bm{\sigma}^{d}\|\|\bm{\sigma}^{f}\|}$
and we obtain the following counterpart of Proposition 4.2.
Proposition 5.2.
The following equality holds, for every $K>0$,
$\overline{C}^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f})=\Phi_{2}(\bar{g}_{1},\bar{g}_{2},\bar{\rho})-\widetilde{K}\Phi_{2}(\bar{d}_{1},\bar{d}_{2},\bar{\rho})$
where $\widetilde{K}=e^{-r^{d}T}K$ and
$\bar{d}_{1}$ $=\frac{-\ln\widetilde{K}-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}{\|\bm{\sigma}^{d}\|\sqrt{T}},\quad\bar{d}_{2}=\frac{-\ln\widetilde{K}+\delta^{q}T-\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}T}{\|\bm{\sigma}^{f}\|\sqrt{T}},$
$\bar{g}_{1}$ $=\frac{-\ln\widetilde{K}+\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}{\|\bm{\sigma}^{d}\|\sqrt{T}},\quad\bar{g}_{2}=\frac{-\ln\widetilde{K}+\delta^{q}T+\bm{\sigma}^{f}(\bm{\sigma}^{d})^{*}T-\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}T}{\|\bm{\sigma}^{f}\|\sqrt{T}}.$
Proof.   We observe that
$\overline{C}^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widehat{S}^{d}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\geq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{d}_{T}-\widetilde{K}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\geq K\}}\big{]}$
$={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{S}^{d}_{T}-\widetilde{K}\big{)}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{d}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\}}\big{]}-\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\big{]}=:\bar{I}_{1}-\widetilde{K}\bar{I}_{2}$
where $\widetilde{S}^{d}_{t}:=e^{-r^{d}t}\widehat{S}^{d}_{t}$ so that $d\widetilde{S}^{d}_{t}=\widetilde{S}^{d}_{t}\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}$ and $\widetilde{K}:=e^{-r^{d}T}K$. The dynamics of $\widehat{S}^{d}$ and $\widehat{S}^{f}$ under ${{\mathbb{Q}}}^{d}$ are
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}r^{d}\,dt+\bm{\sigma}^{d}\,d\textbf{W}_{t}^{d}\big{)},$
$d\widehat{S}^{f}_{t}=\widehat{S}^{f}_{t}\big{(}(r^{d}+\delta^{q})\,dt+\bm{\sigma}^{f}\,d\textbf{W}_{t}^{d}\big{)},$
where $\delta^{q}:=r^{f}-r^{d}-\bm{\sigma}^{f}(\bm{\sigma}^{q})^{*}$. Let ~$\widetilde{{\mathbb{Q}}}^{d}$ be a probability measure introduced in the proof of Proposition 4.2 Then the dynamics of $\widehat{S}^{d}$ and $\widehat{S}^{f}$ under ~$\widetilde{{\mathbb{Q}}}^{d}$ are
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}(r^{d}+\|\bm{\sigma}^{d}\|^{2})\,dt+\bm{\sigma}^{d}\,d\widetilde{\textbf{W}}_{t}^{d}\big{)},$
$d\widehat{S}^{f}_{t}=\widehat{S}^{f}_{t}\big{(}\big{(}r^{d}+\delta^{q}+\bm{\sigma}^{f}(\bm{\sigma}^{d})^{*}\big{)}\,dt+\bm{\sigma}^{f}\,d\widetilde{\textbf{W}}_{t}^{d}\big{)},$
where $\widetilde{\textbf{W}}^{d}$ is a standard Brownian motion under ~$\widetilde{{\mathbb{Q}}}^{d}$. It is now easy to check that
$\bar{I}_{1}=\widetilde{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\big{]}=\Phi_{2}(\bar{g}_{1},\bar{g}_{2},\bar{\rho}),$
$\bar{I}_{2}={{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\big{]}=\Phi_{2}(\bar{d}_{1},\bar{d}_{2},\bar{\rho}),$
where $\bar{g}_{1},\bar{g}_{2},\bar{d}_{1}$ and $\bar{d}_{2}$ are given in the statement of the proposition. $\Box$
The next result is analogous to Proposition 4.3.
Proposition 5.3.
The following equality holds, for every $K>0$,
$\overline{C}^{*}_{0}(\widehat{S}^{f},K,T,\widehat{S}^{d})=\overline{Q}\Big{(}e^{\delta^{q}T}\Phi_{2}(\bar{h}_{1},\bar{h}_{2},\bar{\rho})-\widetilde{K}\Phi_{2}(\bar{d}_{1},\bar{d}_{2},\bar{\rho})\Big{)}$
where $d_{1}$ and $d_{2}$ are given in Proposition 5.2 and
$\bar{h}_{1}=\frac{-\ln\widetilde{K}+\bm{\sigma}^{f}(\bm{\sigma}^{d})^{*}T-\frac{1}{2}\|\bm{\sigma}^{d}\|^{2}T}{\|\bm{\sigma}^{d}\|\sqrt{T}},\quad\bar{h}_{2}=\frac{-\ln\widetilde{K}+\delta^{q}T+\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}T}{\|\bm{\sigma}^{f}\|\sqrt{T}}.$
Proof.   We have
$\overline{C}^{*}_{0}(\widehat{S}^{f},K,T,\widehat{S}^{d})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\overline{Q}\big{(}\widehat{S}^{f}_{T}-K\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\overline{Q}\big{(}\widetilde{S}^{f}_{T}-\widetilde{K}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\geq K\}}\big{]}$
$={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\overline{Q}\big{(}\widetilde{S}^{f}_{T}-\widetilde{K}\big{)}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\}}\big{]}=\overline{Q}\Big{(}{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{f}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\}}\big{]}-\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}>K,\,\widehat{S}^{f}_{T}>K\big{]}\Big{)}$
$=:\overline{Q}\big{(}e^{\delta^{q}T}\widehat{I}_{1}-\widetilde{K}I_{2}\big{)}.$
We introduce an equivalent probability measure ¯$\overline{{\mathbb{Q}}}^{d}$ on $(\Omega,\mathcal{F}_{T})$ by the Radon-Nikodym density
$\frac{d\overline{{\mathbb{Q}}}^{d}}{d{{\mathbb{Q}}}^{d}}:=e^{\bm{\sigma}^{f}\,\textbf{W}_{T}^{d}-\frac{1}{2}\|\bm{\sigma}^{f}\|^{2}T}.$ (11)
Then the process $\overline{\textbf{W}}_{t}^{d},\,t\in[0,T]$ given by the equality $\overline{\textbf{W}}_{t}^{d}=\textbf{W}_{t}^{d}-(\bm{\sigma}^{f})^{*}t$ is a standard Brownian motion under ^$\widehat{{\mathbb{Q}}}^{d}$ and the dynamics of $\widehat{S}^{d}$ and $\widehat{S}^{f,e}$ under ^$\widehat{{\mathbb{Q}}}^{d}$ are
$d\widehat{S}^{d}_{t}=\widehat{S}^{d}_{t}\big{(}r^{d}\,dt+\bm{\sigma}^{f}(\bm{\sigma}^{d})^{*}\,dt+\bm{\sigma}^{d}\,d\overline{\textbf{W}}_{t}^{d}\big{)},$
$d\widehat{S}^{f}_{t}=\widehat{S}^{f}_{t}\big{(}(r^{d}+\delta^{q}+\|\bm{\sigma}^{f}\|^{2})\,dt+\bm{\sigma}^{f}\,d\overline{\textbf{W}}_{t}^{d}\big{)}.$
Hence it is easy to check that the stated expressions are valid. $\Box$
5.4 Superhedging of an Aggregated Quanto Floor EPS
Let us consider a static superhedging strategy for an aggregated quanto floor EPS with $N=1$. We now have that
$\mbox{\rm EPS\,}^{w,q}_{T}=-p_{1}\,\overline{P}_{T}(S^{w,q},k^{l}_{1},T)+p_{1}\overline{P}_{T}(S^{w,q},1,T)-f_{2}\,\overline{C}_{t}(S^{w,q},k^{g}_{1},T)$
and we denote
$P^{*}_{T}(\widehat{S}^{d},K,T,\widehat{S}^{f}):=\big{(}K-\widehat{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\leq K\}}$
and
$P^{*}_{T}(\widehat{S}^{f},K,T,\widehat{S}^{d}):=\overline{Q}\big{(}K-\widehat{S}^{f}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq K\}}.$
The following corollary gives a static superhedging strategy for an aggregated quanto floor EPS.
Corollary 5.3.
The initial cost of a static superhedging hedging strategy $H^{w,e,s,f}$ for an aggregated quanto floor EPS with $N=1$ satisfies
$H^{w,e,s,f}_{0}$ $=w\Big{(}p_{1}P_{0}(\widehat{S}^{d},1,T)-p_{1}\overline{P}^{*}_{0}(\widehat{S}^{d},k^{l}_{1},T)-f_{2}\overline{C}^{*}_{0}(\widehat{S}^{d},k^{g}_{1},T,\widehat{S}^{f})\Big{)}$
$+(1-w)\Big{(}p_{1}\overline{P}_{0}(\widehat{S}^{f},1,T)-p_{1}\overline{P}^{*}_{0}(\widehat{S}^{f},k^{l}_{1},T,\widehat{S}^{d})-f_{2}\overline{C}^{*}_{0}(\widehat{S}^{f},k^{g}_{1},T,\widehat{S}^{d})\Big{)}.$
Finally, we have the following result where $\bar{h}_{1}$ and $\bar{h}_{2}$ are given in Proposition 5.3.
Proposition 5.4.
The following equalities are valid, for every $K>0$,
$\overline{P}^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f})=\widetilde{K}\Phi_{2}(-\bar{d}_{1},-\bar{d}_{2},\bar{\rho})-\Phi_{2}(-\bar{g}_{1},-\bar{g}_{2},\bar{\rho}),$
$\overline{P}^{*}_{0}(\widehat{S}^{f},K,T,\widehat{S}^{d})=\overline{Q}\Big{(}\widetilde{K}\Phi_{2}(-\bar{d}_{1},-\bar{d}_{2},\bar{\rho})-e^{\delta^{q}T}\Phi_{2}(-\bar{h}_{1},-\bar{h}_{2},\bar{\rho})\Big{)}.$
Proof.   We first compute
$\overline{P}^{*}_{0}(\widehat{S}^{d},K,T,\widehat{S}^{f})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}K-\widehat{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\leq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\big{(}\widetilde{K}-\widetilde{S}^{d}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{f}_{T}\leq K\}}\big{]}$
$=\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f}_{T}<K\big{]}-{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{d}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f}_{T}<K\}}\big{]}.$
To obtain the first asserted equality, we use the change of a probability measure from the proof of Proposition 4.2. For the second equality, we observe that
$\overline{P}^{*}_{0}(\widehat{S}^{f},K,T,\widehat{S}^{d})=e^{-r^{d}T}\,{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\overline{Q}\big{(}K-\widehat{S}^{f}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq K\}}\big{]}={{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\overline{Q}\big{(}\widetilde{K}-\widetilde{S}^{f,e}_{T}\big{)}^{+}\mathds{1}_{\{\widehat{S}^{d}_{T}\leq K\}}\big{]}$
$=\overline{Q}\Big{(}\widetilde{K}{{\mathbb{Q}}}^{d}\big{[}\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f}_{T}<K\big{]}-{{\mathbb{E}}}_{{{\mathbb{Q}}}^{d}}\big{[}\widetilde{S}^{f}_{T}\mathds{1}_{\{\widehat{S}^{d}_{T}<K,\,\widehat{S}^{f}_{T}<K\}}\big{]}\Big{)}.$
and we employ the change of a probability measure introduced in the proof of Proposition 5.3. $\Box$
5.5 Approximate Hedging of an Aggregated Quanto Buffer EPS
Let us finally mention that superhedging and approximate strategies can be developed for an aggregated quanto EPS. Since the arguments are analogous to the case of an aggregated effective EPS we only formulate the following result.
Corollary 5.4.
The initial cost of the approximate strategy $H^{w,q,a}$ for an aggregated quanto buffer EPS satisfies
$H^{w,q,a}_{0}=w\widehat{H}^{d}_{0}+(1-w)\widehat{H}^{f,q}_{0}.$
6 Numerical Study for Domestic and Foreign Markets
In numerical studies, the exchange rate, domestic equity and foreign equity are supposed to be governed by the cross-currency market model (3.2), with parameters reported in Table 1. The volatility of the domestic portfolio is set to be 10%, which is consistent with the level of implied volatility index VIX S&P/ASX 200 of the Australian market in 2024, and for the foreign portfolio it is chosen to be 15%, which is the medium value of the implied volatility of the foreign market during the same period, as measured by VIX S&P 500. The exchange rate USD/AUD is taken to be 1.48 (i.e., USD 1.00 costs AUD 1.48), which represent the approximate mean value of the exchange rate in July 2024, and its volatility is assumed to be 9%, which is based on recent data for USD/AUD currency options. Regarding the choice of correlations in our numerical studies, it should be noted that the Australian equity market does not show a strong correlation with U.S. equity market. For instance, as reported in Liu [13], from January 1, 2000, to June 29, 2018, the correlation between the daily returns on the domestic S&P/ASX 200 index and the foreign S&P 500 index was just 0.14. Therefore, we take the correlation between the domestic and foreign equities to be equal to 0.10, which is a conservative value.
The correlation between the exchange rate and the stock market is also a notable phenomenon in the context of pricing EPSs, though the relationship can be complex and thus it requires an extensive expertise in foreign exchange market. For concreteness, we set the correlation between the domestic equity $S^{d}$ (resp., the foreign equity $S^{f}$) and the exchange rate to be 0.05 (resp., -0.05). This means, for example, that a falling AUD (i.e., an increasing exchange rate $Q$) is likely to be linked with in a rising S&P/ASX 200 index (or a falling S&P 500 index), and vice versa. In practice, these correlations can also be inferred from market quotes for traded options on foreign indices and more exotic foreign exchange products.
Recall that the processes $S^{d},S^{f}$ and $Q$ are governed by the stochastic differential equations (3.2). For the standard Brownian motion $\bm{W}^{d}=(W^{d1},W^{d2},W^{d3})$ under ${\mathbb{Q}}^{d}$, we introduce the correlated Brownian motions $Z^{1},Z^{2}$ and $Z^{3}$ by
$Z^{1}_{t}=W^{d1}_{t},\ \,Z^{2}_{t}=\rho_{12}W^{d1}_{t}+\sqrt{1-\rho_{12}^{2}}\,W^{d2}_{t},\ \,Z^{3}_{t}=\alpha_{1}W^{d1}_{t}+\alpha_{2}W^{d2}_{t}+\alpha_{3}W^{d3}_{t}$
where we denote
$\alpha_{1}:=\rho_{13},\quad\alpha_{2}:=\frac{\rho_{23}-\rho_{12}\rho_{13}}{\sqrt{1-\rho_{12}^{2}}},\quad\alpha_{3}:=\sqrt{1-\rho_{13}^{2}-\alpha_{2}^{2}}.$ (12)
Using these equations, we have computed the volatility vectors reported in Table 1; it is easy to check that they are consistent with our assumptions about the volatilities and correlations of processes $S^{d},S^{f}$ and $Q$, that is, $\|\bm{\sigma}^{d}\|=10\%,\|\bm{\sigma}^{f}\|=15\%,\|\bm{\sigma}^{q}\|=9\%$ and $\rho_{12}=0.10,\rho_{13}=0.05,\rho_{23}=-0.05$. Needless to say, in practice the values of volatilities, correlations and the exchange rate should be chosen to reflect the current market data from options market at the time of pricing of an EPS.
Parameter Value
Domestic short-term interest rate $r^{d}$ 4.35%
Foreign short-term interest rate $r^{f}$ 5.25%
Volatility of the domestic equity $\bm{\sigma}^{d}$ $[10.00,0.00,0.00]$
Volatility of the foreign equity $\bm{\sigma}^{f}$ $[1.50,14.93,0.00]$
Volatility of the exchange rate $\bm{\sigma}^{q}$ $[0.45,-0.50,8.98]$
Initial value of the exchange rate $Q_{0}$ 1.48
Table 1: Parameters for cross-currency market model
The graphs (a) and (b) in Figure 1 are for stock prices expressed in their respective currencies. We observe that the trends of the domestic and foreign equity, and the exchange rate are correlated and the volatility of domestic equity is the highest. The graphs (c) and (d) show the tendencies of stock prices and the exchange rate when they are expressed in the domestic currency where, as usual, the domestic price of the foreign equity is calculated by multiplying the foreign stock price by the current level of the exchange rate.
(a) first path
(b) second path
(c) first path in domestic currency
(d) second path in domestic currency
Figure 1: Trends of stock prices and the exchange rate
6.1 Pricing of EPSs on Separated Domestic and Foreign Returns
After defining the parameters of domestic and foreign markets, we first examine the pricing of some standard EPSs on separated domestic and foreign portfolios. We focus here on two particular instances of a standard EPS, the buffer EPS and the floor EPS, and we analyze the hedging cost for an EPS associated with separate hedging strategies, which were discussed in Section 3. We only consider here the case of one year maturity since the influence of maturity is likely to be similar as in the case of EPSs with a single currency, which was extensively studied by Xu et al. [19].
Table 2 presents the hedging cost and key parameters for several variants of domestic and foreign EPSs. For brevity, we use the shorthand notation for various instances of net prices of domestic and foreign EPSs with all prices expressed in the domestic currency for $100$ AUD nominal principal:
(i) Domestic – the price $\mbox{\rm EPS\,}^{d}_{0}$ of the domestic EPS (Corollary 3.2);
(ii) Nominal – the net weighted cost $w\,\mbox{\rm EPS\,}^{d}_{0}+(1-w)Q_{0}\mbox{\rm EPS\,}^{f}_{0}$ where $\mbox{\rm EPS\,}^{f}_{0}$ is the price of the nominal foreign EPS (the foreign market variant of Corollary 3.2);
(iii) Effective – the net weighted cost $w\,\mbox{\rm EPS\,}^{d}_{0}+(1-w)\mbox{\rm EPS\,}^{f,e}_{0}$ where $\mbox{\rm EPS\,}^{f,e}_{0}$ is the price of the effective foreign EPS (Corollary 3.4);
(iv) Quanto – the net weighted cost $w\,\mbox{\rm EPS\,}^{d}_{0}+(1-w)\mbox{\rm EPS\,}^{f,q}_{0}$ where $\mbox{\rm EPS\,}^{f,q}_{0}$ is the price of the quanto EPS (Corollary 3.5).
Buffer $w$ $l_{1}$ $g_{1}$ $p_{2}$ $f_{2}$ Domestic Nominal Effective Quanto
[1] 50% -5% 5% 0.5 0.5 -1.431 -2.231 -1.620 -2.264
[2] 50% -5% 5% 0.8 0.5 -1.169 -1.642 -1.045 -1.675
[3] 20% -5% 10% 0.5 0.5 -0.548 -1.358 -0.789 -1.396
[4] 50% -5% 10% 0.5 0.5 -0.548 -1.055 -0.699 -1.077
[5] 80% -5% 10% 0.5 0.5 -0.548 -0.751 -0.608 -0.760
[6] 20% -5% 10% 0.8 0.5 -0.285 -0.574 -0.027 -0.610
[7] 50% -5% 10% 0.8 0.5 -0.285 -0.466 -0.124 -0.488
[8] 80% -5% 10% 0.8 0.5 -0.285 -0.358 -0.221 -0.367
[9] 50% -5% 10% 0.8 0.8 -0.877 -1.687 -1.118 -1.724
[10] 20% -10% 10% 0.8 0.5 -0.763 -1.633 -0.940 -1.672
[11] 50% -10% 10% 0.8 0.5 -0.763 -1.307 -0.873 -1.331
[12] 80% -10% 10% 0.8 0.5 -0.763 -0.980 -0.807 -0.990
Floor $w$ $l_{1}$ $g_{1}$ $p_{1}$ $f_{2}$ Domestic Nominal Effective Quanto
[1] 50% -5% 5% 0.8 0.5 -0.860 -1.783 -1.353 -1.813
[2] 50% -5% 10% 0.5 0.5 -0.355 -1.142 -0.892 -1.164
[3] 20% -5% 10% 0.8 0.5 0.024 -0.984 -0.706 -1.018
[4] 50% -5% 10% 0.8 0.5 0.024 -0.606 -0.433 -0.627
[5] 80% -5% 10% 0.8 0.5 0.024 -0.228 -0.159 -0.237
[6] 50% -5% 10% 0.8 0.8 -0.568 -1.828 -1.427 -1.863
[7] 50% -10% 10% 0.8 0.5 0.501 0.235 0.317 0.215
[8] 20% -15% 10% 0.5 0.5 0.0504 -0.585 -0.292 -0.619
[9] 50% -15% 10% 0.5 0.5 0.0504 -0.347 -0.164 -0.368
[10] 80% -15% 10% 0.5 0.5 0.0504 -0.109 -0.035 -0.117
[11] 20% -15% 10% 0.8 0.5 0.672 0.662 0.768 0.633
[12] 50% -15% 10% 0.8 0.5 0.672 0.666 0.732 0.648
[13] 80% -15% 10% 0.8 0.5 0.672 0.670 0.696 0.662
[14] 50% -15% 10% 0.8 0.8 0.081 -0.555 -0.262 -0.588
Table 2: Prices of EPSs on separate domestic and foreign returns
From Table 2, the influence of protection and fee legs, and their participation rates are the same for separated cross-currency hedging strategies as for single currency underlying asset, which has been discussed by Xu et al. [19]. Obviously, when we add the weights of domestic and foreign assets into consideration, the hedging costs for this cross-currency EPS will be closer to that for only domestic assets if the weight $w$ of the domestic asset in the underlying portfolio increases, no matter which hedging strategies are used.
With the setting of the predetermined exchange rate equal to the initial exchange rate, we can see that the price of the quanto EPS is close to the price of the nominal foreign EPS, for both buffer and floor EPSs. This can be expected because of the assumption of independence the domestic and foreign assets. Under independence, the prices of standard and quanto European options on a foreign stock are very close. In contrast, we can observe that the assumption of volatility for domestic and foreign assets has a profound effect on the price of the quanto EPS.
Moreover, the prices of effective EPS has the highest values among the three separate cross-currency hedging strategies. This result is under our expectation that only the effective EPS consider the stochastic exchange rate which covers more risk than the other two separate hedging strategies which assume a fixed exchange rate. If the weight $w$ for the domestic asset goes up, then the influence of hedging costs for foreign equities decreases and it will make the difference between three separate cross-currency hedging strategies become smaller.
Example 6.1.
As an illustration of static hedging, we consider the buffer EPS No. 6 from Table 2 with the notional principal AUD 1M with $20\%$ invested in domestic assets and $80\%$ held in foreign assets. Specifically, we assume that the domestic EPS and the effective foreign EPS have the buffer level $l_{1}=-5\%$ with the protection rate $p_{2}=80\%$, the fee level $g_{1}=10\%$ with the fee rate $f_{2}=50\%$, and maturity in one year. The initial prices of assets are not very important here because we focus on their returns. Here we assume that $S^{d}_{0}$ and $S^{f}_{0}$ (roughly) represent market quotes of ASX 200 and S&P 500, $S^{d}_{0}=76.50$ AUD and $S^{f}_{0}=52.50$ USD, and the exchange rate $Q_{0}=1.48$.
For the domestic EPS, one can find from Proposition 3.1 and the Black-Scholes formula that the fair premium equals $20\%\times(-2850)=-570$, which is used to hedge the domestic component of the EPS using call and put options written on $S^{d}$. For the foreign component, from the foreign market variant of Proposition 3.1 we find that the provider can get AUD 216 from building the hedging portfolio using European options written on the process $S^{f,e}$ with $S^{f,e}_{0}=77.7$ AUD. Specifically, the static hedge is composed of:
  • $N^{f,d}p_{2}/S^{f,e}_{0}$ long put options with strike $K^{l}_{1}(S^{f,e}_{0}):=(1+l_{1})S^{f,e}_{0}=73.815$ AUD,
  • $N^{f,d}f_{2}/S^{f,e}_{0}$ short call options with strike $K^{g}_{1}(S^{f,e}_{0}):=(1+g_{1})S^{f,e}_{0}=85.47$ AUD.
We conclude that to statically hedge the foreign component the provider needs to long $(0.8M)\times 0.8/77.7=8237$ units of put options with strike $73.815$ AUD and short $(0.8M)\times 0.5/77.7=5148$ units of call options with strike $85.47$ AUD. The fair premium for the combined position in two EPSs equals $20\%\times(-2850)+80\%\times(-270)=-570-216=-786$ AUD per 1M of the nominal principal.
6.2 Pricing of EPSs on Aggregated Domestic and Foreign Returns
Using once again the data reported in Table 1 we will analyze prices of aggregated effective and quanto EPSs, as given by Corollary 4.1 and Corollary 5.1, respectively. Recall that in both cases we deal with European basket options on domestic returns combined with either effective or quanto foreign returns. Since for closed-form solutions for basket options with arithmetic averaging are not available in lognormal models, we will use instead four numerical approaches, which were studied in Section 4:
(i) Simulation – the classical Monte Carlo approach applied to payoffs of various kinds of EPSs;
(ii) Geometric – the closed-form solutions for basket options with geometric averaging, which were obtained in Propositions 4.1 and 5.1 for effective and quanto EPSs, respectively;
(iii) Moments – the approximate prices obtained by the moment matching method for the aggregated effective EPS from Lemma 4.3 and its variant for the aggregated quanto EPS obtained using Lemma 5.1 instead of Lemma 4.1;
(iv) Super – the superhedging costs $H^{w,e,s,b}_{0}$ and $H^{w,e,s,f}_{0}$ (resp., $H^{w,q,s,b}_{0}$ and $H^{w,q,s,f}_{0}$) from Corollaries 4.2 and 4.3 (resp., Corollaries 5.2 and 5.3).
Buffer $w$ $l_{1}$ $g_{1}$ $p_{2}$ $f_{2}$ Simulation Geometric Moments Super
[1] 50% -5% 5% 0.5 0.5 -1.459 -1.458 -1.476 -0.378
[2] 50% -5% 5% 0.8 0.5 -1.172 -1.170 -1.192 0.197
[3] 20% -5% 10% 0.5 0.5 -0.750 -0.740 -0.770 0.544
[4] 50% -5% 10% 0.5 0.5 -0.580 -0.568 -0.594 0.349
[5] 80% -5% 10% 0.5 0.5 -0.485 -0.489 -0.491 0.155
[6] 20% -5% 10% 0.8 0.5 -0.157 -0.144 -0.182 1.306
[7] 50% -5% 10% 0.8 0.5 -0.292 -0.280 -0.309 0.924
[8] 80% -5% 10% 0.8 0.5 -0.286 -0.294 -0.294 0.542
[9] 50% -5% 10% 0.8 0.8 -0.927 -0.909 -0.950 0.559
[10] 20% -10% 10% 0.8 0.5 -0.986 -0.964 -1.007 0.393
[11] 50% -10% 10% 0.8 0.5 -0.807 -0.790 -0.822 0.174
[12] 80% -10% 10% 0.8 0.5 -0.674 -0.676 -0.680 -0.044
Floor $w$ $l_{1}$ $g_{1}$ $p_{2}$ $f_{2}$ Simulation Geometric Moments Super
[1] 50% -5% 5% 0.8 0.5 -0.882 -0.898 -0.899 1.007
[2] 50% -5% 10% 0.5 0.5 -0.399 -0.398 -0.411 0.856
[3] 20% -5% 10% 0.8 0.5 -0.424 -0.439 -0.443 2.163
[4] 50% -5% 10% 0.8 0.5 -0.003 -0.008 -0.016 1.734
[5] 80% -5% 10% 0.8 0.5 0.106 0.102 0.099 1.305
[6] 50% -5% 10% 0.8 0.8 -0.638 -0.637 -0.657 1.369
[7] 50% -10% 10% 0.8 0.5 0.511 0.502 0.497 2.044
[8] 20% -15% 10% 0.5 0.5 -0.116 -0.130 -0.137 1.376
[9] 50% -15% 10% 0.5 0.5 0.042 0.042 0.028 1.103
[10] 80% -15% 10% 0.5 0.5 0.075 0.068 0.067 0.831
[11] 20% -15% 10% 0.8 0.5 0.855 0.832 0.830 2.638
[12] 50% -15% 10% 0.8 0.5 0.702 0.695 0.685 2.131
[13] 80% -15% 10% 0.8 0.5 0.610 0.597 0.600 1.624
[14] 50% -15% 10% 0.8 0.8 0.068 0.067 0.044 1.765
Table 3: Prices of aggregated effective EPSs under alternative valuation methods
Buffer $w$ $l_{1}$ $g_{1}$ $p_{2}$ $f_{2}$ Simulation Geometric Moments Super
[1] 50% -5% 5% 0.5 0.5 -1.538 -1.539 -1.551 -1.199
[2] 50% -5% 5% 0.8 0.5 -1.336 -1.340 -1.352 -0.609
[3] 20% -5% 10% 0.5 0.5 -0.896 -0.888 -0.913 -0.255
[4] 50% -5% 10% 0.5 0.5 -0.629 -0.623 -0.639 -0.152
[5] 80% -5% 10% 0.5 0.5 -0.513 -0.516 -0.519 -0.049
[6] 20% -5% 10% 0.8 0.5 -0.494 -0.486 -0.515 0.530
[7] 50% -5% 10% 0.8 0.5 -0.427 -0.424 -0.440 0.437
[8] 80% -5% 10% 0.8 0.5 -0.338 -0.343 -0.345 0.344
[9] 50% -5% 10% 0.8 0.8 -1.006 -0.996 -1.022 -0.244
[10] 20% -10% 10% 0.8 0.5 -1.130 -1.113 -1.145 -0.532
[11] 50% -10% 10% 0.8 0.5 -0.816 -0.806 -0.825 -0.406
[12] 80% -10% 10% 0.8 0.5 -0.688 -0.689 -0.692 -0.280
Floor $w$ $l_{1}$ $g_{1}$ $p_{2}$ $f_{2}$ Simulation Geometric Moments Super
[1] 50% -5% 5% 0.8 0.5 -0.969 -0.979 -0.983 0.465
[2] 50% -5% 10% 0.5 0.5 -0.399 -0.397 -0.408 0.548
[3] 20% -5% 10% 0.8 0.5 -0.441 -0.451 -0.457 1.709
[4] 50% -5% 10% 0.8 0.5 -0.060 -0.063 -0.070 1.467
[5] 80% -5% 10% 0.8 0.5 0.106 0.102 0.099 1.224
[6] 50% -5% 10% 0.8 0.8 -0.638 -0.635 -0.653 0.877
[7] 50% -10% 10% 0.8 0.5 0.329 0.319 0.315 1.873
[8] 20% -15% 10% 0.5 0.5 -0.282 -0.290 -0.301 1.017
[9] 50% -15% 10% 0.5 0.5 -0.081 -0.084 -0.093 0.878
[10] 80% -15% 10% 0.5 0.5 0.020 0.014 0.013 0.740
[11] 20% -15% 10% 0.8 0.5 0.489 0.471 0.464 2.420
[12] 50% -15% 10% 0.8 0.5 0.449 0.439 0.433 1.995
[13] 80% -15% 10% 0.8 0.5 0.515 0.505 0.506 1.570
[14] 50% -15% 10% 0.8 0.8 -0.129 -0.134 -0.149 1.405
Table 4: Prices of aggregated quanto EPSs under alternative valuation methods
In Table 3 (resp., Table 4) we report prices of aggregated effective EPSs (resp., aggregate quanto EPSs) for $100$ AUD nominal principal, which were obtained using four alternative numerical approaches. As a remark, the Monte Carlo estimates are based on 1000000 simulations of stock prices and thus the prices of cross-currency basket options, and hence also the prices of cross-currency EPSs, obtained through simulation are considered to be exact values. It is clear that the three numerical approaches for aggregated hedging costs - simulation, geometric averaging, and three moment matching - provide very close results, with no more than 0.01% difference in most cases. However, we also note that, when compared with three moment matching method, the geometric average approximation provides prices closer to the exact prices (Simulation) in more situations. Moreover, the aggregated hedging costs under three moment matching method are usually cheaper then the exact costs, no matter for buffer or floor EPS. As was expected, the superhedging prices are always higher than the exact prices since the terminal value of a superhedging strategy dominates the payoff of an EPS and the market model is arbitrage-free.
It should be noted that the prices of aggregated effective EPSs and aggregated quanto EPSs with identical protection and fee parameters are very close. This is due to the fact that we use a quanto exchange rate $\overline{Q}=Q_{0}$ and the volatility for the exchange rate is relatively small (as compared to stocks), so that the influence of the exchange rate is also relatively small. In reality, a holder may choose a predetermined exchange rate different from its current value to satisfy their requirements and then the prices of aggregated effective EPSs and aggregated quanto EPSs would be markedly different. It is worth noting that in Tables 3 and 4 the prices of aggregated effective EPSs are higher than that of aggregated quanto EPSs when dealing with both buffer EPSs and floor EPSs, because the protection of effective EPS which contains the influence of stochastic exchange rate is more comprehensive than quanto EPS.
Comparing aggregated hedging and superhedging strategies, it is clear that a aggregated superhedging cost (reported as Super in Tables 3 and 4) is higher than the costs of a respective aggregated effective hedging strategy and quanto hedging strategy in all situations, considering all three numerical approaches for aggregated hedging costs. This result is consistent with the definitions and structures of hedging strategies: the superhedging strategy covers more than the overall potential gain or loss and thus it costs more than the respective hedging strategy. No matter whether the hedging cost is positive (EPS is more beneficial to investors) or negative (EPS is more beneficial to providers), only the provider needs to do hedging and thus he will pay more under joint superhedging strategies than aggregated hedging strategies.
Moreover, the structure of EPS products will influence the difference between aggregated hedging costs and superhedging costs. We can see that the difference between hedging and superhedging costs for floor EPS are higher than for buffer EPS, under both effective EPS and quanto EPS. Because the structure of floor EPS is more complex than buffer EPS, it needs more long and short positions of different European options to hedge the final payoff under floor EPS and it increases the costs rapidly when applying superhedging strategies.
Example 6.2.
As an example of an aggregated effective EPS, we consider the floor EPS No. 5 from Table 3 with the notional principal AUD 1M split into 0.8M invested in domestic assets and 0.2M held in foreign assets. The buffer level equals $l_{1}=-5\%$ with participation rate $p_{2}=80\%$, the fee level $g_{1}=10\%$ with participation rate $f_{2}=50\%$, and the maturity date $T=1$. As shown in Section 5, the arbitrage-free premium for the EPS equals AUD 1060.
We first note that $S^{w,e}_{0}=0.8\times 76.5+0.2\times 1.48\times 52.5=76.74$ AUD. Hence a static hedging portfolio with European options written on $S^{w,e}$ involves:
  • $\frac{p_{2}}{S^{w,e}_{0}}$ short put options with strike $K^{l}_{1}(S^{w,e}_{0}):=(1+l_{1})S^{w,e}_{0}=72.903$,
  • $\frac{p_{2}}{S^{w,e}_{0}}$ long put options with strike $S^{w,e}_{0}=76.74$,
  • $\frac{f_{2}}{S^{w,e}_{0}}$ short call options with strike $K^{g}_{1}(S^{w,e}_{0}):=(1+g_{1})S^{w,e}_{0}=84.414$.
The provider needs to short $(0.8M/76.74)=10425$ units of put options with strike 72.903, long $(0.8M/76.74)=10425$ units of put options with strike 76.74, and short $(0.5M/76.74)=6516$ units of call options with strike 84.414.
7 Conclusions
In this paper, the studies of equity protection swaps (EPS) from Xu et al. [19] have been expanded into a more general situation - with consideration of cross-currency reference portfolio. An EPS is an insurance product for the variable annuity that provides protection against potential losses and collects insurance fees from actual gains. Compared with guarantees attached to variable annuities, which are offered in the U.S. market, an EPS has a simple structure and, more importantly, is a standalone financial derivative, which can be purchased independently of the variable annuity itself. We contend that it is a suitable product for the superannuation market in Australia where a typical holder of a pension account is able to directly invest in domestic and foreign securities, with 3.5 trillion Australian dollars (AUD) of total assets currently held in superannuation accounts. It should be stressed that an EPS can be structured in such a way that its initial cost is null and a fee is only charged when a reference portfolio generates significant profits during the lifetime of an EPS.
According to Xu et al. [19], the EPS products in a single economy can be perfectly hedged by a static portfolio of European call and put options. We have shown here that it is still possible to use European options to hedge the cross-currency EPS product but the form of a static hedging should be adapted to an EPS at hand and the use of basket options may be required for some cross-currency products. We first examine an EPS with separated protections for domestic and foreign portfolios, including nominal, effective, and quanto foreign returns. Then the basic algorithm is similar to the standard EPS product in a single currency, as presented in Xu et al. [19]. A comparison of hedging costs for EPSs with three types of foreign equity returns shows that the hedging cost for foreign equity’s effective return is the highest one, which was expected since an EPS of that kind provides the most comprehensive protection against currency and equity risk. Moreover, when a quanto exchange rate is chosen to be equal to the initial exchange rate, one expects that the hedging cost for nominal foreign return and quanto return should be close to each other and this is indeed the case, as can be seen from numerical results. Therefore, if a buyer of an EPS has strong views about the future level of the exchange rate, then the EPS referencing the quanto return could be a suitable product.
Another important class of cross-currency products are EPSs referencing the aggregated return on bespoke cross-currency portfolios. We examine the case of the cross-currency effective return, as well as the quanto return, and we use cross-currency basket options to hedge EPS products on aggregated returns. To this end, we present three different methods to pricing the basket option that is needed for EPS product valuation, namely, the geometric averaging method, the moments matching technique, and the Monte-Carlo simulation. We provide a numerical study to compare the approximate basket option prices with simulation results, which are considered to be exact values. It can be seen that the hedging costs for cross-currency EPS products are fairly close under three valuation methods. Since cross-currency basket options on bespoke portfolios are not likely to be easily available in practice, we also examine superhedging strategies based solely on the same domestic and foreign market options as in the case of separated protections.
Finally, we also provide a superhedging strategy for the cross-currency EPS product under consideration of aggregated returns. By evaluating the structure of the cross-currency EPS product and analysing its aggregated returns, we can split the aggregated returns into two parts and build a superhedging strategy, which can be built by using exotic European options with a single currency equity as underlying assets. We can see that superhedging costs are higher than other aggregated hedging costs for all EPS and this superhedging strategy has costs heavily related to the EPS structures that the costs for EPS products with more complex structures are higher. We may draw a conclusion that the EPS providers’ hedging strategies should depend on the EPS buyers’ needs that determine the properties of the EPS products. In the meantime, the quanto exchange rate should be applied if the EPS buyer has any assumptions of the future exchange rate.